### v1

# GAD-Generative Learning for HD Map-Free Autonomous Driving

Abstract
Deep-learning-based techniques have been widely adopted for autonomous driving software stacks for mass production in recent years, focusing primarily on perception modules, with some work extending this method to prediction modules. However, the downstream planning and control modules are still designed with hefty handcrafted rules, dominated by optimization-based methods such as quadratic programming or model predictive control. This results in a performance bottleneck for autonomous driving systems in that corner cases simply cannot be solved by enumerating hand-crafted rules. We present a deep-learning-based approach that brings prediction, decision, and planning modules together with the attempt to overcome the rule-based methods’ deficiency in real-world applications of autonomous driving, especially for urban scenes. The DNN model we proposed is solely trained with 10 hours of human driver data, and it supports all mass-production ADAS features available on the market to date. This method is deployed onto a Jiyue test car with no modification to its factory-ready sensor set and compute platform. the feasibility, usability, and commercial potential are demonstrated in this article.
1 Introduction
Autonomous driving software on the market to date usually is designed following a classical engineering methodology called “divide and conquer”, where the problem is divided into smaller sub-tasks to reduce complexity. These sub-tasks usually include perception, prediction, motion planning, and control. However, this easily interpretable rule-based method is highly dependent on accurate prior knowledge, which severely restricts its scalability. High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information, including topology at both lane and road levels, lane geometries, crosswalks, traffic lights, intersections, etc. This prior information, with superior accuracy, greatly lessens the burden of perception and enormously facilitates planning tasks for autonomous driving. Unfortunately, HD maps are proven hard to scale up due to the complexity and costliness of map generation and maintenance, as well as failure to reflect changes in real-world road conditions timely, which is often constrained by resource-intensive collection processes. For instance, lane reduction or expansion, changes in merge or diverge points, and lane closure or lane shifts due to roadblocks, barricades, and temporary construction zones are often seen in reality promptly but reflected on HD maps with months’ of delay.
Furthermore, the rule-based planning module is held accountable for most failures and deficiencies of the autonomous driving system in real driving scenarios, especially for urban scenes, hence it is often regarded as the bottleneck of the entire software stack. New rules must be designed and re-tuned in accordance with different driving scenarios and new corner cases. This process is expensive and it scales poorly for the simple fact that the scenario space for the SDV is infinite while rules can only be of limited number.
End-to-end autonomous driving aims to consider sub-tasks in a full-stack manner jointly have been proposed to resolve the above issues. Methods[3, 21, 22, 24, 48, 49, 38, 7, 25, 40, 20] that direct learning from raw sensor data to planning trajectories or driving commands eliminate the need for manual planning rules and prerequisite HD maps, leading to efficient data utilization and adaptability to diverse driving scenarios. However, those works lack interpretability since omitting modularized results and are difficult to optimize. Interpretability is of key importance in a safety-critical system, particularly if a bad event were to happen in mass production. Moreover, the absence of modularized structure and prior knowledge makes these methods very brittle when deployed in real-world closed-loop.
To overcome the aforementioned problems, we explore an alternative method which runs entirely on the edge and purely relied on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [12] based on perception results. The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first. We believe this is effective for two primary reasons. Firstly, landmarks(lane center, lane divider, road boundary, and pedestrian crossing) provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized. Second, for turning scenes (left turn, right turn, u-turn), we gather routing information from SD map or historical commuting trajectory and avoid collisions with the help of real-time occupancy.
To break the bottleneck brought by rule-based planning but maintain the system’s explainability and stability, we take modularized outputs from perception and propose a general framework to handle the prediction-planning tasks in a data-driven and integrated manner. Our framework consists of a trajectory generator and evaluator. The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric info in grid form to validate and score candidates with physically feasible output that can navigate in HD map-free. We validate our framework in a closed loop by deploying and testing it in real-world user-level autonomous driving vehicles.
We illustrate the pros of our design that decouple perception from the end-to-end stack: Firstly, the decoupled tasks are both constructed in an end-to-end way, benefiting from efficient data utilization but maintaining system’s stability and robustness. Secondly, the prediction-planning integration omits the parameterized prediction output, modeled in prior Gaussian distribution, which cannot always hold, thus avoiding information loss for downstream planning. The perception task is deterministic based, as long as the measurement is sufficient, therefore, less affected. Thirdly, the original tools for perception and planning can be reused for these two tasks. Otherwise, prerequisites of 3d construction at the sensor level like Nerf[14] are heavily resource-consuming.
The real-world closed-loop validation needs to be highlighted as well for the following reasons: Firstly, a closed-loop planer will adjust to the target in the loop to keep safe and comfortable since real-world planning needs to take into account the deviations brought by the controller’s modeling errors and external disturbances, which are omitted and thus cannot be tested by the open-loop planner. Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller. The open-loop tests end when the trajectories are generated and compared with the recorded ones. However, unexpected and unreasonable jerks and curvatures, which can only be tested when the controller is in the loop, are usually ignored, and thus, the results become less convincing. This leaves the effectiveness of those methods (like UniAD[22], VAD[24], mile[20], ST-P3[21]) unknown for onboard road tests and mass production.
Our key contributions are summarized as follows:
  1. 1.
    We propose an industry-level paradigm for data-driven prediction-planning with an HD map-free setting. Our model leverages non-parametric information with better tolerance to upstream perception and handling prediction and planning tasks in one backbone.
  2. 2.
    To the best of our knowledge, we are the first evaluation of such system in a complex, real-world, urban road with a factory-ready sensor set and compute platform. We demonstrate the imperative of closed-loop evaluation compared to imitation metric adopted by extensive offline studies.
  3. 3.
    We extend the state-of-the-art data-driven approach and propose a method that goes beyond simple imitation or trajectory sampling and instead reasons about the potential in generative learning by combining max marge planning with multi-modal imitation.
2 Related work
2.1 End to End Learning
End-to-end autonomous driving and its effectiveness have been investigated recently. There exists a body of research [40, 47, 20, 24] focused on closed-loop end-to-end driving within simulators [10, 27]. However, a domain gap persists between the simulator environment and the real world, particularly concerning sensor data and the motion status of agents. Recently, open-loop end-to-end autonomous driving has attracted more attention. End-to-end autonomous driving methods [22, 24] that involve learning intermediate tasks claim their effectiveness in improving final planning performance. Nevertheless, the effectiveness of open-loop demonstration has been challenged recently[28], in terms of imbalanced data distribution and metric limitation. We tackle and bypass that controversial issue by validating our method in a real-world closed-loop road test.
Our work combines actor perception and scene occupancy as inputs for our paradigm. First, we demonstrate how off-the-shelf planners can use parametric results to enhance the quality of trajectory generation. Second, we show how to evaluate the generated trajectory with scene information as additional supervision. Third, we demonstrate the feasibility of combining scene-level occupancy and parametric landmarks as a replacement for an HD map in an urban area.
2.2 Data Driven Prediction
For actor-based trajectory predictions, these methods either generate a fixed set of trajectories [4, 2, 32, 50], or draw samples via generative modeling to characterize the distribution [36, 41, 17, 26]. In order to encode scene context, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8]. More recently, methods have started to leverage the graph obtained from HD-map in order to better represent lane connectivity. VectorNet [13] encodes map and agent as polylines with a global interaction graph. LaneGCN [29] treats actor past and the lane graph separately and then fuses them with a series of attention layers. Temporal occupancy maps [23, 37, 19], on the other hand, provide another option to forecast grid or heat map at the scene level. We follow the philosophy of [11, 18, 38, 42, 46] in modeling scene-level representations, but propose a multi-stage pipeline to perform forecasting in both formats. Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are mainly adopted in the safety layer and to boost some of our samplers.
2.3 Data Driven Planning
For implicit planning methods, the network directly generates planned trajectory or control commands with direct and simple design[33, 20, 1, 5, 25, 30]. But it suffers from robustness and lack of interpretability[6]. For trajectory-based methods, modeling strategies from trajectory prediction are adopted, which build an imitation decoder trained under the supervision of human demonstration. PlanT[35] proposed a transformer-based backbone in actor-level representation and validated in simulation. SafetyNet[44] developed an ML planner and tested it in downtown San Francisco, but it relies on a fallback layer to handle algorithm failure with a traditional rule-based system. VAD[24] models the driving scene as a fully vectorized representation. More recently, explicit methods have been shown to adapt better to more challenging environments. Such approaches usually build a cost map with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost. The cost map is organized in a non-parametric form learned directly from the network[48, 38, 3, 49, 7]. Start with the pioneering work on NMP [48], extended by DSDNet[49] and P3[38] that combined hand-crafted and learning-based costs to obtain an integrated cost volume. In the follow up works, MP3 [3] introduced mapless, while STP3[21] explored vision-based inputs. We also adopt this combination to choose the best trajectory but extend the system’s robustness by boosting the generating algorithm with modularized perception output. In contrast to all the above methods, which solely rely on a rule-based sampler, our paradigm leverages imitation learning to build a diversified hybrid trajectory set and stays safe by being evaluated in an interpretable cost volume as a summary of scene occupancy and agent motion.
Figure 1: Model Structure of GAD. The model initiates with two streams for rasterization and vectorization respectively to encode scene and instance-level features. The rasterization branch utilize convolution layers to derive the scene embedding, which is then decoded into cost volumes for scoring candidate trajectories from multiple samplers and prediction grid maps for agent prediction. Meanwhile the scene embedding are cropped and incorporated with the instance encoding from vectorization branch to decode agents’ future movements.
3 Methods
3.1 Input Representation
We make the assumption that we have access to a black-box perception module which can effectively map low-level sensor data to 3D-tracked entities, visible landmarks, and scene occupancy. At each timestamp, we are provided with measured states for each tracked entity, including their 2D positions, velocities, and accelerations. Furthermore, the perception module also supplies information on all visible landmarks and road boundaries in front of our ego. Stop lines are also provided when they become visible as the SDV approaches the junction area. The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form. The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its movability status, characterized as non-parametric To encode upstream output for different usages, we design representations for scene and instance levels, which are applied to trajectory evaluation and generation respectively.
Scene Representation
We adopt rasterization proposed by[9] for scene representation. Scene rasterization with convolutional forwarding can better capture scene-level cues compared to instance vectorization, avoiding information loss to the maximum extent. Owing to the convolutional network’s property of exploiting spatially local correlation, undefined or irregular elements like construction areas, traffic cones, and roadblocks can be incorporated with a unified design regardless of how they are organized and whether they have regularized shapes. Additional scene information depicted by free space or occupancy, lacking parametric description, can hardly be involved with vectorized representation and, once omitted, can be fatal to safety-critical systems. To serve as a qualified trajectory evaluator, BEV rasterization involves environment cues from online perception outputs as much as possible to capture those ”devils in detail”. We follow the interpretable design as that of in [3], but rasterize different scene elements(e.g. agent, lanes, occupancy, route) into different binary channels to enable reasoning about the distinct element separately. The details can be found in our appendix.
Instance Representation
The instance representation is employed to address the shortcomings of scene encoding, which are afflicted by the truth that, rasterization is prone to distorting highly organized and structured information, rendering it disorganized. Since this procedure is irreversible, making it challenging to retrieve instance encoding for lane center lines, road boundaries, and other elements. The vectorized representation combined with the transformer has been proven to be superior to CNN forwarding on BEV raster in terms of accuracy and efficiency[13] for the motion prediction task. To take advantage of modeling in such form, we adopt vectorization[13] to encode the perception outputs into vector sets. Instances with parametric description are encoded in the vector form, including route points, ego and agent historical, lane center lines, and road boundaries, while non-parametric outputs are ignored, like occupancy, construction area, roadblocks, and other irregular elements. Each element includes position and type features, with the position described relative to the SDV pose. The relationships between separated elements are organized as a graph.
3.2 Model Structure
We build a multi-resolution, two-stream backbone network to encode inputs and extract agents and scene features separately. As shown in Fig. 1. One steam processes the rasterized scene while the other deals with the vectorized graph. As for scene encoding, we build a convolutional encoder composed of several residual blocks with corresponding layers and strides. Namely, the features after each block are down-sampled accordingly from the input in the BEV raster to obtain the scene embedding. The rasterization branch has more layers than the vectorization branch as the input is much denser than the vectorized graph. This stream is responsible for aggregating geometric information from the scene context to extract global appearance and motion cues. After obtaining scene embedding from the CNN encoder, it is passed through two separate decoders with identical structures but different purposes. Both decoders are composed of deconvolution layers to restore the output to the same size as the rasterization input. The first decoder outputs the cost volume while the other outputs the predicted grid map. Both will be illustrated in the next section.
Meanwhile, in the other stream, inspired by [13], the instance-level encoding is built on a hierarchical graph network-based architecture. It consists of a PointNet-based [34] local sub-graph for processing local information from vectorized inputs and a global graph using a Transformer encoder [43] for reasoning about interactions over agents and landmark features.
Finally, the corresponding scene encoding for the agent is obtained via oriented crop[4] in ego-centric mid-level feature and fused with instance encoding by concatenation along the feature dimension which outputs high-dimensional tensors $f_{agent}$. With the agent-level embeddings, an imitation deocder is utilized to output the predicted trajectories per agent. The details for the encoder and decoders are illustrated in the appendix.
3.3 Route
When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the SDV should follow. In HD map-free settings, however, this is not possible, especially for turning scenarios in intersections where no leading landmark is available. In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues. The lane center lines constructed by landmarks are mainly used in cruise and lane change scenarios, while in turning scenarios, the starting and ending positions are recorded in advance, once vehicles approach the intersection, the human commuting trajectories will serve as route. More details can be found in the appendix.
3.4 Prediction
We follow a multi-stage design for scene and agent prediction. In the first stage, the pipeline aims to predict the scene level grid map, also known as the heat map or occupancy. We regress a grid map $G_{ij}^{t}=dec(f_{ras})$ with a resolution matches that of the rasterization input by decoding the middle layer features.
Figure 2: Multi-stage Prediction
The decoder, as mentioned in 3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box. This grid is trained with pixel-wise focal loss inspired from [15]:
$\begin{array}[]{r}L_{occ\_pred}=-\sum_{ij}[\tilde{g_{ij}}(1-\hat{g_{ij}})^{2}log(\hat{g_{ij}})+\\ (1-\tilde{g_{ij}})\hat{g_{ij}}^{2}log(1-\hat{g_{ij}})]\end{array}$ (1)
Here, $\tilde{g_{ij}}$ represents the class label for the grid, where $\tilde{g_{ij}}=1$ indicates that it is occupied.
In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid. For the agent-level prediction, we follow the interactive design. Since the output trajectories are mainly adopted in sampling for SDV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego. In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior. Those landmarks indicate areas where interaction is most likely to happen. The agents’ trajectories will be decoded conditioned on that lane prior:
$Traj=dec(f_{agent}|f_{lane})$ (2)
Here, $f_{agent}$ is the instance encoding as illustrated in Sec. 3.2, $f_{lane}$ is the corresponding lane encoding obtained in the same way. For the training phase, the ground truth $\tilde{lane}$ corresponding to the agent’s future trajectory will be labeled and incorporated as a condition prior:
$L_{traj\_pred}=L2(\tilde{Traj},dec(f_{agent}|f_{\tilde{lane}}))$ (3)
The training loss is measured by the $L2$ distance between the predicted and ground truth trajectory.
3.5 Trajectory Evaluation
The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories. We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost maps for scoring candidate trajectories. We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map. Our evaluator decodes a space-time cost map $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster. The cost of a trajectory is defined as the sum of costs at its space-time way-points. The best candidate trajectory is the one with the lowest cost according to the learned cost:
$s=arg\min_{\hat{s}\in S}\sum_{t}[C(\hat{s})+\alpha O_{cc}(\hat{s})+\beta G_{rid}(\hat{s})]$ (4)
where $S$ represents the set of viable future trajectories, $O_{cc}$ represents perception occupancy, $G_{rid}$ is the predicted grid map, $\alpha$ and $\beta$ are the weights respectively. Since selecting the minimum-cost trajectory within a discrete set is not differentiable, we use the max-margin loss to penalize trajectories that have small cost and are different from the human driving trajectory or are unsafe. Let $\hat{s}$ and $\tilde{s}$ be the input and human trajectory respectively for a given example. We utilize the max-margin loss to encourage the human driving trajectory to have a smaller cost than other trajectories.
$L_{cost\_volume}=\max_{\hat{s}}[C(\tilde{s})-C(\hat{s})+Diff(\tilde{s},\hat{s})]$ (5)
The $Diff$ value here serves as compensation to quantitatively measure the difference between the human demonstrated and our candidates. It has to tell how goodness of the human driving or how badness of our samplers, quantified by planning targets, avoiding static collisions, and managing dynamic interactions, etc. So that the evaluator can assign comparable costs to the trajectories in our candidate set. In particularly:
$Diff(\tilde{s},\hat{s})=L2(\tilde{s},\hat{s})+O(\hat{s})+G(\hat{s})$ (6)
where $O$ is the occupancy cost, $G$ is the rest of the prediction costs as defined in Sec. 3.4. The imitation loss $L2$ measures the $l2$ distance between trajectory $\hat{s}$ and the ground-truth $\tilde{s}$ for the entire horizon.
3.6 Trajectory Generation
Given the learned cost volume, the final trajectory can then be computed by minimizing Eq. 4. Note, however, that this optimization with physical constraints is NP-hard. On one hand, we evaluate and extend previous works on trajectory sampling[48, 3, 38, 21] with a more sophisticated design by incorporating prior knowledge and structured outputs from perception and prediction. Furthermore, we leverage a GAN-based pipeline during the generator’s training to solve the optimization problem. In this section, we describe our sampling and generation algorithm.
The sampling and generating algorithms can be mainly categorized into reference line-based and reference line-free. Methods incorporating reference information can generate long-term consistent outputs in the loops. Since this information rarely changes over time, bridging consecutive frames promotes cross-frame stability. However, scenarios still need to be handled where reference-related information can hardly be obtained. To enhance the scalability of our system, we consider trajectory from the following methods. In our implementation, curve and retrieval-based samplers are reference line-free, while the others rely on reference information constructed by visual landmarks or commuting trajectories.
Curve Based Sampler
As the most basic sampler, it has been adopted by a few previous methods[48]. We follow [48] with little modifications for comfort concerns. We take into account the initial states of a Self-Driving Vehicle(SDV), such as velocity, acceleration, heading, and road curvature where the SDV is located, and then sample a combination of straight lines, circles, and clothoid curves as trajectories with the given initial states and random end states within a pre-defined range. The curve-based sampler serves as the baseline sampling algorithm in our experiment. The details can be found in the appendix.
Retrieval Based Sampler
Following prior work that uses data-driven trajectory sampling techniques in mapless driving scenarios [3], we create an expert trajectory dataset by clustering manual driving data. Once this dataset is constructed during pre-processing, SDV can retrieve trajectories with high similarity according to its current motion states. The construction and retrieval details can be found in the appendix.
Lattice Sampler
Qi added The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the SDV. When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process. The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in longitude. In our paradigm, they are both available as depicted by Sec. 3.4 Sec. 3.1. In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39].Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude. We illustrate the detail in the appendix.
Imitation Planner
For the imitation planner, previous approaches choose to mimic a human driver with one single trajectory[22, 24]. But such designs suffer from model collapse in both longitude and lateral. Owing to our generator-evaluator paradigm, we can generate a bunch of diverse candidates and pass them to the downstream evaluation task. We follow multiple-trajectory[8] design in trajectory prediction to estimate the conditional distribution $P(S|f_{plan})$ in motion planning task. Where $S=[\hat{s_{1}},...\hat{s_{m}}]$ is the trajectories with $m$ different maneuvers such as lane following, lane changing, agent following, agent overtaking. $f_{plan}$ stands for feature prepared for planning task. ||||| | stands for concatenation.
$f_{plan}=[f_{scene}||f_{route}||f_{SDV}]$ (7)
$P(S|f_{plan})=\sum_{M}P(S|m,f_{plan})P(m|f_{plan})$ (8)
Our imitation generator decodes a multitude of $M$ possible trajectories for ego move in local coordinates.
${[\hat{s_{1}},...\hat{s_{m}}]}=G(f_{plan})$ (9)
During the training phase, we then identify mode $m^{\ast}$ that is closest to the ground-truth trajectory according to distance function,
$L_{imit\_plan}=arg\min_{i\in{1,...m}}dist(\tilde{s},\hat{s_{i}})$ (10)
We select $L2$ dist for the distance function.
GAN Based Planner
The above imitation generator is trained separately from our evaluator by different headers. With the attempt to generate trajectory with minimal cost, we propose a GAN base planner with modification to the training paradigm that bridges the isolated optimization targets, so that Eq. 4 can be solved by learning. For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize below function while the discriminator tries to maximize it.
$L_{gan}=\min_{G}\max_{D}(E_{x}[log(D(x))]+E_{z}[log(1-D(G(z)))])$ (11)
$log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly. In our paradigm, we replace the discriminator with an evaluator. But different from the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator output score to measure the discrepancy between candidates quantitatively. So the optimization target for evaluator Eq. 5 still holds, but the learning target of the imitation planner becomes:
$\ L_{generator}=E(G(m))+Diff(\tilde{s},G(m))$ (12)
$L_{evaluator}=\max_{\hat{s}}[E(\tilde{s})-E(\hat{s})+Diff(\tilde{s},\hat{s})]$ (13)
with:
$\ E(s)=C(s)=\sum_{(x_{t},y_{t})\in{s}}C(x_{t},y_{t})$ (14)
$\ C_{xy}^{t}=dec(f_{ras})$ (15)
$\ G(m)=\hat{S}=[\hat{s_{1}},...\hat{s_{m}}]$ (16)
Here m is an hidden variable represent the manual in planning.
3.7 Safety layer
For on-road autonomous driving motion planning, a strategy for safety coverage is necessary. The perception occupancy and trajectory prediction results are used to validate the trajectory candidates with respect to safety. The trajectory prepared for the safety layer is a little bit different to Eq. 2 in that no prior is assumed in decoding. We sort the trajectories according to the evaluated costs and iterate them to check the basic rule in modularized form(like trajectory and bounding box) until we cache enough candidates by ascending order. To ensure cross-frame consistency, we reserve $n$ validated trajectories with the lowest cost and compare them with planning output in the last frame. The above objective promotes the system’s robustness with respect to the corner cases that our model fails to handle. We believe such a layer of safety guarantee is indispensable for manufacturing our system at scale.
4 Experiments
We construct different baseline variants of our models in both open-loop and closed-loop for ablative studies and compare them with state-of-the-art alternative methods quantitatively. We gather in-house datasets for training and open-loop settings. It contains a wide range of driving scenarios in a densely populated urban environment. We collected the record files in 15 hours, which captured the perception output, the SDV trajectory, and route info. All the records are post-processed to obtain 200k clips in total. By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future. We partition the dataset into 16w for training and 4w for testing. To validate the potential of our method, we disable the functionality in the safety layer as well as the optimizer and safety backup subsystem in other methods. More details on protocols are provided in the appendix.
4.1 Prediction
We focus on the $L2$ distance and hit rate for the open-loop prediction. Preview SOTAs are evaluated in our HD map-free setting.
4.2 Open-loop Planning
For the open-loop planning, we focus on two evaluation metrics: $L2$ error and collision rate, and align the planning horizon to 3.0s for a fair comparison. We compute the $L2$ error between the planned trajectory and the human driving trajectory, and evaluate how often the SDV would collide with other agents on the road. Detailed comparison with previous methods is shown in Tab. Note that the Vanilla algorithm is penalized based on how much the trajectory deviates from the ground truth, thus it achieves the lowest L2 error but largest collision rates in all prediction horizons.
4.3 Closed-loop Experimental
We first evaluate the necessity of different sampler on closed-loop simulation by ablation study. Then, find the best combination to evaluate in the real-world. The results of ablation study illustrated appendix.
We extensively tested andn appendix. compare the above methods in the real world, in densely populated urban area, under the supervision of human safety drivers. The driving route used for our experiments is shown on appendix. During the 100+ mile public road testing, the model successfully performed a wide variety of challenging maneuvers including lane-following, merging, yielding to pedestrians or nudging around parked cars.
5 Conclusions
6 Acknowledgements
We would like to thank the following contributors who work on the toolchains to support our closed-loop evaluation and onboard road test: Chi.Zhang, Yue.Li, Bolin.Zhao, Yanze.Liu, Liang.Zhu.
Appendix A Appendix
A.1 Scene Rasterization
The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections. Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect. To distinct SDV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history. The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness. we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route. Since we explore route from different source, we cast different routing form into one single uniformed *** All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
A.2 Instance Vectorization
A.3 Model Structure
A.4 Route
As indicated previously, the route plays a critical role in guiding SDVs along roadways. Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the SDV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes. Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions. Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the SDV to change lanes.
Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the SDV lacking appropriate instructions to navigate intersections. To address this, the starting and ending positions for turning are recorded. As the SDV approaches intersections, human-driving trajectories are employed as the route to guide the SDV through the intersections.
A.5 Curve Based Sampler
In this appendix, we present the sampling procedure with more details. Vehicle trajectories can be categorized into two primary types: straight and turning. And turning trajectories encompass lane changes and turning at intersections. Lane-changing trajectories are characterized by variable curvature over time, while turning at intersections follows circular-like curves. As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.
For straight ones, a sequence of acceleration values are sampled to account for various velocity-changing scenarios, which are subsequently integrated over time to obtain speed profiles. With initial states(the position and heading angle) of the autonomous vehicle, a set of straight trajectories can be obtained.
Circle trajectories employ the bicycle model[31] to establish the relationship between steering angle $\phi$ and path curvature $\kappa$, represented by the formula $\kappa=2tan(\phi)/L\approx 2\phi/L$, where $L$ is the distance between the front and rear axles of the autonomous vehicle. Using a constant curvature and different sampled speed profiles, a set of circle trajectories can be derived.
Clothoid curves involve a proportional relationship between the curvature $\kappa$ of a point and its distance to the original point $\xi$, which can be defined as
$\mathbf{s(\xi)}=\mathbf{s_{0}}+a\left[C\left(\frac{\xi}{a}\right)\mathbf{T_{0}}+S\left(\frac{\xi}{a}\right)\mathbf{N_{0}}\right]$ (1)
[…]
Where $\mathbf{s(\xi)}$ is the clothoid curve, $\mathbf{T_{0}}$ and $\mathbf{N_{0}}$ are the tangent and normal vector of point $\mathbf{s_{0}}$, $S(\xi)$ and $C(\xi)$ are associated coefficients and defined as the Fresnel integral. As the shapes of clothoid curves solely determined by the scaling factor a, we can generate a variety of clothoid trajectories by considering a suitable range for a, in combination with the initial curvature obtained from the bicycle model and the speed profiles.
A.6 Lattice Sampler
The lattice sampler is a reference line-based method used to acquire multiple candidate trajectories along the reference line, taking into account future movements of surrounding obstacles to avoid potential hazards.
To begin, this method involves sampling multiple candidate trajectories based on the Frenet Frame, where the smooth lane central line is projected to the S coordinate and its normal direction is the L coordinate. Trajectories are composed of longitudinal and lateral paths, with the longitudinal one described as the station and time, denoted as $S=\left\{s(t)\right\}$, and the lateral path described as the lateral station and longitudinal one denoted, denoted $L={l(s)}$.
When sampling trajectories, given the initial state $[s_{0},s_{0}^{{}^{\prime}},s_{0}^{{}^{\prime\prime}}],[l_{0},l_{0}^{{}^{\prime}},l_{0}^{{}^{\prime\prime}}]$ at time $t_{0}$, the end states $[s_{1},s_{1}^{{}^{\prime}},s_{1}^{{}^{\prime\prime}}],[l_{1},l_{1}^{{}^{\prime}},l_{1}^{{}^{\prime\prime}}]$ at time $t_{1}$ are sampled to cover cruising, following, overtaking and stopping maneuvers, and consequently fits polynomials.
As for cruising maneuvers, vehicles must adhere to a specific cruising speed by adjusting their acceleration and deceleration. Therefore a set of prediction time horizon $t_{1}$ and end-condition longitudinal velocities $s_{1}^{{}^{\prime}}$ are uniformly sampled within a pre-defined range of the cruising speed. This sampling process is facilitated by a quartic polynomial.
When it comes to stopping maneuvers, vehicles are obliged to come to a complete stop within a designated distance in order to comply with traffic signals or ensure a safe distance from stationary obstacles. In this scenario, the end states such as speed and accelerations are fixed at 0, while the prediction time $t_{1}$ and end longitudinal distance $s_{1}$ are sampled. These parameters are then described using a quintic polynomial.
Regarding following and overtaking maneuvers, these actions often arise when a preceding vehicle is moving at a slower pace or a new vehicle unexpectedly merges into a lane. The execution of these complex maneuvers is facilitated through the use of an s-t graph, which depicts the occupation of the lane by vehicles over time as illustrated in Figure x. By analyzing the s-t graph, end states can be sampled within unoccupied areas, marked in red (for overtaking) and blue (for following), thus facilitating the completion of these maneuvers through the use of a quintic polynomial.
Finally, these polynomials can be transformed back to Cartesian coordinate system to obtain the corresponding trajectories, which are consequently evaluated for compliance with kinematic constraints. If any trajectory does not meet the requirements, it will be removed from the sampling set.
A.7 Retrieval-based Sampler
The retrieval based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the SDV’s current state from a pre-established expert trajectory dataset.
To build the expert trajectory dataset, we utilized about 350 hours of driving trajectories from experienced human drivers. However, a substantial portion of the original human driving trajectories contains duplicate or very similar samples. Drawing inspiration from [3], we conducted subdivision and clustering on the original human driving trajectories in order to preserve diversity while simplifying the expert trajectory dataset.
To subdivide the original trajectories into different bins, we use the initial velocity v, acceleration a and curvature $\kappa$, with respective bin sizes of $1.0\frac{m}{s}$, $0.5\frac{m}{s^{2}}$, $0.01\frac{1}{m}$. In each bins, trajectories are clustered into 2000 sets and the closest trajectory to each cluster center is retained. Finally, the trajectories from each subset will be aggregated to the comprehensive expert trajectory dataset.
[…]
$\textit{D}=\beta_{1}\cdot|v-v_{SDV}|+\beta_{2}\cdot|a-a_{SDV}|+\beta_{3}\cdot|\kappa-\kappa_{SDV}|$ (1)
In this paper, we assign different weights to the differences in velocity v, acceleration a and curvature $\kappa$, with $\beta_{1}=1.0$, $\beta_{2}=5.0$, $\beta_{3}=40.0$ and all nodes with $D<1.0$ will be added into output for the retrieval-based sampler.
Figure 3: pipeline
A.8 Hardware
Our autonomous driving algorithms have undergone rigorous testing on the recently launched Jiyue 01 vehicle, whose detailed parameters can be found in the following link: https://www.jiyue-auto.com/robocar-01. The Jiyue 01 is equipped with a comprehensive sensor suite, which includes 11 cameras strategically positioned to provide front, side, rear, and surround-view coverage. Additionally, the vehicle features 5 millimeter-wave radars, enabling accurate perception in challenging weather conditions, and 12 ultrasonic radars for close-range detection.
Moreover, the Jiyue 01 is fitted with 2 NVIDIA Orin-X chips, representing a powerful and advanced autonomous driving hardware and software platform. Specifically engineered for autonomous driving, these chips offer a processing capability of 508 TOPS, meeting the high demands of algorithm design and ensuring optimal performance. More NVIDIA-Orin parameters can be found in the following link: https://developer.nvidia.com/drive/agx.
A.9 Task Perf
Figure 4: Test Drive Route
A.10 Road Test
Our road tests are conducted in the Jiading District of Shanghai, China, known for ”Shanghai Automobile City”. The detailed testing route is illustrated in Figure 4, encompassing straight stretches, lane changes, intersections, and turning scenes involving various traffic participants.
The route begins at No.1688 Yecheng Road, a road with two lanes in each direction, encounters frequent cars and electric bicycles. It passes through two intersections controlled by traffic lights before entering a three-lane structure and making a careful lane change to the right-most lane, where sometimes occupied by static vans. It then queues before the stop line and then turns onto Shengxin Road, a four-lane-per-direction road heavily congested commuter route with significant traffic flows in all directions.
While weaving through vehicles on Shengxin Road and go straight across the first intersection, the route carefully enters the right-most lane, and then turns right, encountering electric bicycles and vehicles from different directions at the intersection. After turning onto South Huicheng Road, a one lane per direction road often occupied by passengers and tricycles, the traffic flows are slowed down, requiring careful navigations.
The route then follows a right-turning at the next intersection and enters Huocheng Road, a narrow one-lane-per-direction road where trucks from the opposite direction sometimes encroaching on the ego lane, necessitating slight steering to avoid collisions. Proceeding through the next intersection, the route goes straight and executes an unprotected left turn to cross a bustling transportation route frequently used by trucks. Finally, the route circles back to Yecheng Road and returns to destination No.1688.
A.11 Model structure
Our model architecture consists of two main branches. The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding. Subsequently, the output from the rasterization branch is fed into two identical-structured decoders, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch. The decoders generate cost volumes and prediction grid maps.
Concurrently, the vectorization branch processes parametric instance information through a PointNet subgraph, which includes three edge convolutions. The output from this branch is then passed through a downstream Transformer encoder to obtain the instance encoding.
Following the extraction of the scene embedding, features by oriented crop around the agents are obtained. These features are then processed through a sequence of maxpooling and a multi-layer perceptron (MLP). The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
A.12 Training details
The Adam optimizer is used with an initial learning rate of $5xe^{-4}$ and a batch size of 32. The scheduler of reduce on loss plateau decay is also implemented, with a reducing factor of 0.3 and a patience factor of 2, which means that if there is no improvement after 2 epochs, the learning rate will be reduced by 0.3 times the original value. The lower bound of the learning rate is set to $10^{-8}$ and the model parameters are randomly initialized. Experiments are conducted with 4 NVIDIA H800 GPUs with 60 epochs to complete.
Add two stage training @weijian.sun

### v2

@@
 We present a deep-learning-based approach that brings prediction, decision, and planning modules together with the attempt to overcome the rule-based methods’ deficiency in real-world applications of autonomous driving, especially for urban scenes.
-The DNN model we proposed is solely trained with 10 hours of human driver data, and it supports all mass-production ADAS features available on the market to date.
+The DNN model we proposed is solely trained with 10 hours of valid human driver data, and it supports all mass-production ADAS features available on the market to date.
 This method is deployed onto a Jiyue test car with no modification to its factory-ready sensor set and compute platform.
 The feasibility, usability, and commercial potential are demonstrated in this article.
+[^1] Test records will be available at https://github.com/weijian-sun/GAD
 1 Introduction
-Autonomous driving software on the market to date usually is designed following a classical engineering methodology called “divide and conquer”, where the problem is divided into smaller sub-tasks to reduce complexity.
+Autonomous driving software on the market to date is usually designed following a classical engineering methodology called “divide and conquer”, where the problem is divided into smaller sub-tasks to reduce complexity.
 These sub-tasks usually include perception, prediction, motion planning, and control.
 However, this easily interpretable rule-based method is highly dependent on accurate prior knowledge, which severely restricts its scalability.
-High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information, including topology at both lane and road levels, lane geometries, crosswalks, traffic lights, intersections, etc.
-This prior information, with superior accuracy, greatly lessens the burden of perception and enormously facilitates planning tasks for autonomous driving.
+High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information, greatly lessens the burden of perception and enormously facilitates planning tasks for autonomous driving.
 Unfortunately, HD maps are proven hard to scale up due to the complexity and costliness of map generation and maintenance, as well as failure to reflect changes in real-world road conditions timely, which is often constrained by resource-intensive collection processes.
-For instance, lane reduction or expansion, changes in merge or diverge points, and lane closure or lane shifts due to roadblocks, barricades, and temporary construction zones are often seen in reality promptly but reflected on HD maps with months’ of delay.
-Furthermore, the rule-based planning module is held accountable for most failures and deficiencies of the autonomous driving system in real driving scenarios, especially for urban scenes, hence it is often regarded as the bottleneck of the entire software stack.
+Furthermore, the rule-based planning module is held accountable for most failures and deficiencies of the autonomous driving system in real driving scenarios, especially for urban scenes.
+Hence, it is often regarded as the bottleneck of the entire software stack.
 New rules must be designed and re-tuned in accordance with different driving scenarios and new corner cases.
-This process is expensive and it scales poorly for the simple fact that the scenario space for the SDV is infinite while rules can only be of limited number.
-End-to-end autonomous driving aims to consider sub-tasks in a full-stack manner jointly have been proposed to resolve the above issues.
-Methods[3, 21, 22, 24, 48, 49, 38, 7, 25, 40, 20] that direct learning from raw sensor data to planning trajectories or driving commands eliminate the need for manual planning rules and prerequisite HD maps, leading to efficient data utilization and adaptability to diverse driving scenarios.
-However, those works lack interpretability since omitting modularized results and are difficult to optimize.
+This process is expensive and it scales poorly for the simple fact that the scenario space for the Autonomous Drive Vehicle(ADV) is infinite while rules can only be of limited number.
+Figure 1: Comparison on the design of model-based planning.
+(a)Imitation planner with single trajectory output ([22, 24, 35, 44]).
+(b)Sample a bunch of candidates and evaluate with cost volume([48, 3, 38, 21]).
+(c)Proposed generative method, in which the data-driven generator gets feedback from the evaluator.
+End-to-end autonomous driving, which aims to consider sub-tasks in a full-stack manner jointly, has been proposed to resolve the above issues.
+Methods[22, 24, 25, 40, 20, 33, 5, 30] that direct learning from raw sensor data to planning trajectories or driving commands eliminate the need for hand-crafted heuristic rules and the prerequisite HD maps, avoiding infomation loss and better adapt to diverse scenarios.
+However, these approaches lack interpretability which plays a key role in understanding and tracing how the system envolves, thus will be difficult for the developer to optimize.
 Interpretability is of key importance in a safety-critical system, particularly if a bad event were to happen in mass production.
 Moreover, the absence of modularized structure and prior knowledge makes these methods very brittle when deployed in real-world closed-loop.
-To overcome the aforementioned problems, we explore an alternative method which runs entirely on the edge and purely relied on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [12] based on perception results.
-The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first.
+To overcome the aforementioned problems, we explore an alternative method that runs entirely on the edge and purely relies on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [12] based on perception results.
+The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first, primarily due to considerations of scalability and mass production feasibility.
 We believe this is effective for two primary reasons.
-Firstly, landmarks(lane center, lane divider, road boundary, and pedestrian crossing) provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized.
-Second, for turning scenes (left turn, right turn, u-turn), we gather routing information from SD map or historical commuting trajectory and avoid collisions with the help of real-time occupancy.
-To break the bottleneck brought by rule-based planning but maintain the system’s explainability and stability, we take modularized outputs from perception and propose a general framework to handle the prediction-planning tasks in a data-driven and integrated manner.
+Firstly, landmarks provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized.
+Secondly, for turning scenes, we gather routing information from historical commuting trajectories and avoid collisions with the help of real-time occupancy.
+However, the proposed methodology is not limited to commuting scenarios; it supports a versatile selection of route formats.
+We also validated its non-commuting variant in the experiments.
+To break the bottleneck brought by rule-based planning but maintain the system’s explainability and stability, we propose a general framework to handle the prediction-planning tasks in a data-driven and integrated manner.
 Our framework consists of a trajectory generator and evaluator.
-The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric info in grid form to validate and score candidates with physically feasible output that can navigate in HD map-free.
+The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric info in grid form to validate and score candidates [^2] We note that by parametric form we mean the structured numerical representations of certain perception outputs, such as the observed vehicle states.
+Conversely, the non-parametric form refers to unstructured representations that cannot be described by parameters, such as occupancy..
 We validate our framework in a closed loop by deploying and testing it in real-world user-level autonomous driving vehicles.
-We illustrate the pros of our design that decouple perception from the end-to-end stack: Firstly, the decoupled tasks are both constructed in an end-to-end way, benefiting from efficient data utilization but maintaining system’s stability and robustness.
-Secondly, the prediction-planning integration omits the parameterized prediction output, modeled in prior Gaussian distribution, which cannot always hold, thus avoiding information loss for downstream planning.
-The perception task is deterministic based, as long as the measurement is sufficient, therefore, less affected.
-Thirdly, the original tools for perception and planning can be reused for these two tasks.
-Otherwise, prerequisites of 3d construction at the sensor level like Nerf[14] are heavily resource-consuming.
-The real-world closed-loop validation needs to be highlighted as well for the following reasons: Firstly, a closed-loop planer will adjust to the target in the loop to keep safe and comfortable since real-world planning needs to take into account the deviations brought by the controller’s modeling errors and external disturbances, which are omitted and thus cannot be tested by the open-loop planner.
+We elucidate the rationale behind our architectural that segregates perception from the comprehensive end-to-end stack: Firstly, for mass production friendly concern, the modularized perception provides explaninablity and tunablitiy, Secondly, information articulated in parametric form facilitates reasoning at the agent level, not solely for application within the safety layer and the analysis of explainable cases, but also to enhance our trajectory generation processes.
+Thirdly, parameterization encompasses post-tracking and fusion for the results of single or multi-modal detection on raw sensor data which remain formidable unresolved issues, particularly when considering the substitution of these processes with a model to attain comparable results.
+We leave perception involved version to follow up works when those issue resolved in the future.
+The real-world closed-loop validation needs to be highlighted as well: Firstly, a on road planer will adjust to the target in the loop to keep safe and comfortable since real-world planning needs handle the deviations brought by the external disturbances and controller’s modeling errors, which is hard to simuate and omitted by open-loop planner as well.
 Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller.
-The open-loop tests end when the trajectories are generated and compared with the recorded ones.
-However, unexpected and unreasonable jerks and curvatures, which can only be tested when the controller is in the loop, are usually ignored, and thus, the results become less convincing.
-This leaves the effectiveness of those methods (like UniAD[22], VAD[24], mile[20], ST-P3[21]) unknown for onboard road tests and mass production.
+Since unexpected and unreasonable jerks and curvatures can only be tested when the controller is in the loop with the planner.
+The above limitations make the results in open-loop or simulator less convincing and leaves the effectiveness of those methods (like UniAD[22], VAD[24], mile[20], ST-P3[21]) unknown for onboard road tests and mass production.
 Our key contributions are summarized as follows:
@@
     We extend the state-of-the-art data-driven approach and propose a method that goes beyond simple imitation or trajectory sampling and instead reasons about the potential in generative learning by combining max marge planning with multi-modal imitation.
+As depicted in Fig.
+1
 2 Related work
-2.1 End to End Learning
-End-to-end autonomous driving and its effectiveness have been investigated recently.
-There exists a body of research [40, 47, 20, 24] focused on closed-loop end-to-end driving within simulators [10, 27].
-However, a domain gap persists between the simulator environment and the real world, particularly concerning sensor data and the motion status of agents.
-Recently, open-loop end-to-end autonomous driving has attracted more attention.
-End-to-end autonomous driving methods [22, 24] that involve learning intermediate tasks claim their effectiveness in improving final planning performance.
-Nevertheless, the effectiveness of open-loop demonstration has been challenged recently[28], in terms of imbalanced data distribution and metric limitation.
-We tackle and bypass that controversial issue by validating our method in a real-world closed-loop road test.
-Our work combines actor perception and scene occupancy as inputs for our paradigm.
-First, we demonstrate how off-the-shelf planners can use parametric results to enhance the quality of trajectory generation.
-Second, we show how to evaluate the generated trajectory with scene information as additional supervision.
-Third, we demonstrate the feasibility of combining scene-level occupancy and parametric landmarks as a replacement for an HD map in an urban area.
+2.1 End to End Learning&Validation
+End-to-end autonomous driving has become a significant research focus, with numerous recent studies.
+These methods, using neural networks, produce planning trajectories or control commands, with many[40, 47, 20] focusing on closed-loop simulator experiments [10, 27].
+However, a persistent domain discrepancy exists between the simulated environment and the actual world, especially in terms of sensor data fidelity and the dynamic states of agents.
+The accurate modeling of agent behaviors and the simulation of real-world data[14] within closed-loop environments continue to pose formidable challenges that have yet to be resolved.
+Recently, open-loop experiments with real-world data have attracted more attention[24, 22, 21, 48].
+Methods like UniAD[22] claim their effectiveness in improving final planning performance by involving learning intermediate tasks.
+However, it relies on an additional optimal control module to obtain a feasible trajectory and prevent collisions, which violates the initial purpose of breaking optimization-based deficiencies in motion planning task.
+Nevertheless, the effectiveness of open-loop demonstration has been challenged by [28] in terms of imbalanced data distribution and metric limitation.
+We address these contentious issues by rigorously validating our approach through real-world closed-loop road test, wherein the proposed methodology eschews the dependence on conventional optimization algorithms to derive trajectories that adhere to specified constraints.
 2.2 Data Driven Prediction
-For actor-based trajectory predictions, these methods either generate a fixed set of trajectories [4, 2, 32, 50], or draw samples via generative modeling to characterize the distribution [36, 41, 17, 26].
+Prior research has extensively explored actor-based prediction methods for future driving dynamics.
+These methods either produce specific trajectories [4, 2, 32, 51] or use generative models to outline probable outcomes [36, 41, 17, 26].
 In order to encode scene context, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8].
-More recently, methods have started to leverage the graph obtained from HD-map in order to better represent lane connectivity.
+More recent methods have started to leverage the topologic graph obtained from HD maps in order to better represent lane connectivity.
 VectorNet [13] encodes map and agent as polylines with a global interaction graph.
@@
 Temporal occupancy maps [23, 37, 19], on the other hand, provide another option to forecast grid or heat map at the scene level.
-We follow the philosophy of [11, 18, 38, 42, 46] in modeling scene-level representations, but propose a multi-stage pipeline to perform forecasting in both formats.
+We follow the philosophy of [11, 18, 38, 42, 46] in modeling scene-level representations but propose a multi-stage pipeline to perform forecasting in both formats.
 Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are mainly adopted in the safety layer and to boost some of our samplers.
 2.3 Data Driven Planning
-For implicit planning methods, the network directly generates planned trajectory or control commands with direct and simple design[33, 20, 1, 5, 25, 30].
-But it suffers from robustness and lack of interpretability[6].
-For trajectory-based methods, modeling strategies from trajectory prediction are adopted, which build an imitation decoder trained under the supervision of human demonstration.
+For imitation based methods, modeling strategies from trajectory prediction are adopted, which build an imitation decoder trained under the supervision of human demonstration.
+Apart from the aformention end to end works, SafetyNet[44] developed a machine learning planner and tested it in downtown San Francisco, but the model can hardly avoid collision without the help of a fallback layer.
+VAD[24] modeled the driving scene as a fully vectorized representation.
 PlanT[35] proposed a transformer-based backbone in actor-level representation and validated in simulation.
-SafetyNet[44] developed an ML planner and tested it in downtown San Francisco, but it relies on a fallback layer to handle algorithm failure with a traditional rule-based system.
-VAD[24] models the driving scene as a fully vectorized representation.
-More recently, explicit methods have been shown to adapt better to more challenging environments.
-Such approaches usually build a cost map with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost.
-The cost map is organized in a non-parametric form learned directly from the network[48, 38, 3, 49, 7].
-Start with the pioneering work on NMP [48], extended by DSDNet[49] and P3[38] that combined hand-crafted and learning-based costs to obtain an integrated cost volume.
-In the follow up works, MP3 [3] introduced mapless, while STP3[21] explored vision-based inputs.
+However, the monotonous single trajectory imitation suffers from robustness and lacks interpretability[6], thus difficult to tune by case and hard to scale up.
+More recently, explicit methods with great interpretability have been shown to adapt better to more challenging environments.
+Such approaches usually build a cost volume with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost.
+The cost volume is organized in a non-parametric form learned directly from the network[48, 38, 3, 49, 7].
+Beginning with pioneering work on NMP [48], DSDNet[49] and P3[38] made further developments that combined hand-made and learning-based costs to obtain an integrated cost volume.
+In the follow-up work, MP3 [3] introduced a mapless approach, while ST-P3[21] explored vision-based inputs.
 We also adopt this combination to choose the best trajectory but extend the system’s robustness by boosting the generating algorithm with modularized perception output.
-In contrast to all the above methods, which solely rely on a rule-based sampler, our paradigm leverages imitation learning to build a diversified hybrid trajectory set and stays safe by being evaluated in an interpretable cost volume as a summary of scene occupancy and agent motion.
+In contrast to all methods above, which solely rely on a rule-based sampler, our model leverages multi-modal imitation learning to build a diversified hybrid trajectory set and stays safe by being evaluated in an interpretable cost volume as a summary of scene occupancy and agent motion.
 Figure 2: Model Structure of GAD.
 The model initiates with two streams for rasterization and vectorization respectively to encode scene and instance-level features.
-The rasterization branch utilize convolution layers to derive the scene embedding, which is then decoded into cost volumes for scoring candidate trajectories from multiple samplers and prediction grid maps for agent prediction.
-Meanwhile the scene embedding are cropped and incorporated with the instance encoding from vectorization branch to decode agents’ future movements.
+The rasterization branch utilizes convolution layers to derive the scene embedding, which is then decoded into cost volumes for scoring candidate trajectories from multiple samplers and prediction grid maps for agent prediction.
+Meanwhile the scene embedding are cropped and incorporated with the instance encoding from vectorization branch to decode ego trajectory and agents’ future movements.
 3 Methods
 3.1 Input Representation
-We make the assumption that we have access to a black-box perception module which can effectively map low-level sensor data to 3D-tracked entities, visible landmarks, and scene occupancy.
-At each timestamp, we are provided with measured states for each tracked entity, including their 2D positions, velocities, and accelerations.
-Furthermore, the perception module also supplies information on all visible landmarks and road boundaries in front of our ego.
-Stop lines are also provided when they become visible as the SDV approaches the junction area.
-The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form.
-The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its movability status, characterized as non-parametric To encode upstream output for different usages, we design representations for scene and instance levels, which are applied to trajectory evaluation and generation respectively.
+The goal of this paper is to present a deep learning (DL) pipeline that goes from prediction to motion planning in an integrated manner; hence, it is not our intention to cover perception at length.
+In our work, the perception module follows a typical mainstream design that provides tracked entities, landmarks, and occupancy voxels (also known as volumn pixels) by incorporating images from a suite of independent cameras with different field of views (FoVs) to formulate a surround view at 360 degrees.
+The perception module has a deep neural net that directly stitches features in bird’s-eye view (BEV) form, and transformers are used to capture both spatial and temporal contexts.
+Occupancy is also an integral part of outputs by this perception module, providing a useful scene representation of a three-dimensional world that offers free space for SDV to navigate through.
+More details are provided in the appendix.
+To encode upstream output for different usages, we design representations for scene and instance levels [^3] We note that by scene we mean the overall environment context for the current frame, by instance we mean the elements in the scene such as ego, agent, landmark, route, etc., which are applied to trajectory evaluation and generation, respectively.
 Scene Representation
@@
 Additional scene information depicted by free space or occupancy, lacking parametric description, can hardly be involved with vectorized representation and, once omitted, can be fatal to safety-critical systems.
-To serve as a qualified trajectory evaluator, BEV rasterization involves environment cues from online perception outputs as much as possible to capture those ”devils in detail”.
+To serve as a qualified trajectory evaluator, BEV rasterization involves environment cues from online perception outputs as much as possible to capture those “devils in detail”.
 We follow the interpretable design as that of in [3], but rasterize different scene elements(e.g.
@@
 Instances with parametric description are encoded in the vector form, including route points, ego and agent historical, lane center lines, and road boundaries, while non-parametric outputs are ignored, like occupancy, construction area, roadblocks, and other irregular elements.
-Each element includes position and type features, with the position described relative to the SDV pose.
+Each element includes position and type features, with the position described relative to the ADV pose.
 The relationships between separated elements are organized as a graph.
 3.2 Model Structure
-We build a multi-resolution, two-stream backbone network to encode inputs and extract agents and scene features separately.
+We build a two-stream backbone network to encode inputs and extract agents and scene features separately.
 As shown in Fig.
-1.
+2.
 One steam processes the rasterized scene while the other deals with the vectorized graph.
-As for scene encoding, we build a convolutional encoder composed of several residual blocks with corresponding layers and strides.
-Namely, the features after each block are down-sampled accordingly from the input in the BEV raster to obtain the scene embedding.
-The rasterization branch has more layers than the vectorization branch as the input is much denser than the vectorized graph.
-This stream is responsible for aggregating geometric information from the scene context to extract global appearance and motion cues.
-After obtaining scene embedding from the CNN encoder, it is passed through two separate decoders with identical structures but different purposes.
-Both decoders are composed of deconvolution layers to restore the output to the same size as the rasterization input.
-The first decoder outputs the cost volume while the other outputs the predicted grid map.
-Both will be illustrated in the next section.
+To elaborate, the rasterization stream employs a convolutional encoder of multiple resolutions to derive the scene embedding, responsible for aggregating geometric information from the scene context to extract global appearance and motion cues.
+The scene embedding is then passed through two separate deconvolution decoders to generate the cost volumes and the predicted grid maps respectively.
 Meanwhile, in the other stream, inspired by [13], the instance-level encoding is built on a hierarchical graph network-based architecture.
 It consists of a PointNet-based [34] local sub-graph for processing local information from vectorized inputs and a global graph using a Transformer encoder [43] for reasoning about interactions over agents and landmark features.
-Finally, the corresponding scene encoding for the agent is obtained via oriented crop[4] in ego-centric mid-level feature and fused with instance encoding by concatenation along the feature dimension which outputs high-dimensional tensors $f_{agent}$.
-With the agent-level embeddings, an imitation deocder is utilized to output the predicted trajectories per agent.
-The details for the encoder and decoders are illustrated in the appendix.
+Finally, the corresponding scene encoding for the agent is derived via oriented crop[4] in ego-centric scene feature and fused with instance encoding to obtain agent-level features $f_{agent}$.
+Then an imitation decoder is utilized to produce the trajectories for agents and the ego.
+The details for the encoders and decoders are illustrated in the appendix.
 3.3 Route
-When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the SDV should follow.
-In HD map-free settings, however, this is not possible, especially for turning scenarios in intersections where no leading landmark is available.
+When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the ADV should follow.
+However, in HD map-free settings, this is not possible, especially for turning scenarios in intersections where no leading landmark is available.
 In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues.
@@
 Figure 3: Multi-stage Prediction
-The decoder, as mentioned in 3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box.
+The decoder, as mentioned in Sec.
+3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box.
 This grid is trained with pixel-wise focal loss inspired from [15]:
@@
 For the agent-level prediction, we follow the interactive design.
-Since the output trajectories are mainly adopted in sampling for SDV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego.
+Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego.
 In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior.
@@
 The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories.
-We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost maps for scoring candidate trajectories.
+We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost volumes for scoring candidate trajectories.
 We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map.
-Our evaluator decodes a space-time cost map $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster.
+Our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster.
 The cost of a trajectory is defined as the sum of costs at its space-time way-points.
@@
 However, scenarios still need to be handled where reference-related information can hardly be obtained.
-To enhance the scalability of our system, we consider trajectory from the following methods.
-In our implementation, curve and retrieval-based samplers are reference line-free, while the others rely on reference information constructed by visual landmarks or commuting trajectories.
-Curve Based Sampler
-As the most basic sampler, it has been adopted by a few previous methods[48].
-We follow [48] with little modifications for comfort concerns.
-We take into account the initial states of a Self-Driving Vehicle(SDV), such as velocity, acceleration, heading, and road curvature where the SDV is located, and then sample a combination of straight lines, circles, and clothoid curves as trajectories with the given initial states and random end states within a pre-defined range.
-The curve-based sampler serves as the baseline sampling algorithm in our experiment.
-The details can be found in the appendix.
-Retrieval Based Sampler
-Following prior work that uses data-driven trajectory sampling techniques in mapless driving scenarios [3], we create an expert trajectory dataset by clustering manual driving data.
-Once this dataset is constructed during pre-processing, SDV can retrieve trajectories with high similarity according to its current motion states.
-The construction and retrieval details can be found in the appendix.
+To enhance the scalability of our system, we follow [48, 3] to build curve and retrieval-based sampler in reference-line free and illustrate them in the appendix.
+The proposed lattice sampler, imitation planner, and GAN-based planner will be explained in this section.
+They incorperate reference info constructed by visual landmarks or commuting trajectories.
 Lattice Sampler
-Qi added The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the SDV.
+The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the ADV.
 When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process.
@@
 3.1.
-In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39].Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude.
+In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39].
+Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude.
 We illustrate the detail in the appendix.
@@
 ||||| | stands for concatenation.
-$f_{plan}=[f_{scene}||f_{route}||f_{SDV}]$ (7)
+$f_{plan}=[f_{scene}||f_{route}||f_{ADV}]$ (7)
 $P(S|f_{plan})=\sum_{M}P(S|m,f_{plan})P(m|f_{plan})$ (8)
@@
 4 can be solved by learning.
-For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize below function while the discriminator tries to maximize it.
+For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize the following function while the discriminator tries to maximize it.
 $L_{gan}=\min_{G}\max_{D}(E_{x}[log(D(x))]+E_{z}[log(1-D(G(z)))])$ (11)
@@
 In our paradigm, we replace the discriminator with an evaluator.
-But different from the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator output score to measure the discrepancy between candidates quantitatively.
+But unlike the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator outputs a score to measure the discrepancy between candidates quantitatively.
 So the optimization target for evaluator Eq.
@@
 $\ G(m)=\hat{S}=[\hat{s_{1}},...\hat{s_{m}}]$ (16)
-Here m is an hidden variable represent the manual in planning.
+Here, m is a hidden variable representing the manual in planning.
+However, the trajectory evaluation Eq.
+14 involve index trajectory points in the learned cost volume $C_{xy}^{t}=dec(f_{ras})$, which is non-differentiable, thus cannot be back propagated to the generator.
+To address this issue, we use a soft function to get a meaningful gradient.
+The grid_sample operator[1] is adopted here to obtain the gradients via interpolation between values in our cost volume.
 3.7 Safety layer
@@
 4 Experiments
-We construct different baseline variants of our models in both open-loop and closed-loop for ablative studies and compare them with state-of-the-art alternative methods quantitatively.
-We gather in-house datasets for training and open-loop settings.
-It contains a wide range of driving scenarios in a densely populated urban environment.
-We collected the record files in 15 hours, which captured the perception output, the SDV trajectory, and route info.
-All the records are post-processed to obtain 200k clips in total.
-By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future.
-We partition the dataset into 16w for training and 4w for testing.
-To validate the potential of our method, we disable the functionality in the safety layer as well as the optimizer and safety backup subsystem in other methods.
-More details on protocols are provided in the appendix.
+We develop various baseline variants of our models under both open-loop and closed-loop configurations for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies.
+We gather in-house datasets for both training and open-loop evaluations, encompassing an extensive array of driving scenarios within densely populated urban settings.
+To substantiate the feasibility of the data-driven planners, we deactivate the functionalities in the safety layer as well as the optimizer and safety backup subsystems in comparative methodologies such as optimal control in UniAD[22], etc.
 4.1 Prediction
-We focus on the $L2$ distance and hit rate for the open-loop prediction. Preview SOTAs are evaluated in our HD map-free setting.
+For the prediction task, prior state-of-the-art methodologies are appraised within HD map-free context, wherein solely the observable landmarks are furnished to the predictors.
+We measure the performance of methods on root mean squared error and hit-rate overall future timesteps $t\in{1,2,3}$ seconds.
+In Tab.
+1, we compare all methods using the same inputs.
+In general, our multi-modal design with interactive prior outperformed our reimplemented TNT[50] with 16 candidates, while the uni-modal version shared similar results with the current state-of-the-art.
+Method Mean L2 / hit-rate $<$ 1m Avg.
+1s 2s 3s
+VectorNet[13] 0.75/0.80 1.72/0.40 2.85/0.22 1.34
+GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35
+TNT[50] 0.65/0.78 1.01/0.57 1.72/0.34 0.90
+GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
+Table 1: Prediction Results in uni-modal and multi-modal.
+Method trajectory sources L2(m) Collision(%)
+curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s
+UniAD[22] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78
+VAD[24] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69
+NMP[48] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05
+MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97
+P3[38] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04
+GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52
+GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48
+GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07
+GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
+Table 2: Open loop planning results of imitation based method, sampling based method and GAD.
+Method comfort event in 30Km efficiency MPI
+Lat. accel. Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute
+UniAD[22] 0.351 3.12 3.68 - - - - 15.20 0.31
+ST-P3[21] 0.857 4.54 8.53 4 3 14 3 8.62 1.36
+GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72
+GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75
+GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75
+GAD-GAN 0.511 1.15 3.52 3 1 2 1 5.51 4.28
+DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
+Table 3: Road Test Outcomes: Data were systematically collected during a 30 km road test, encompassing the computation of average time per loop.
+Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded.
+The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
 4.2 Open-loop Planning
-For the open-loop planning, we focus on two evaluation metrics: $L2$ error and collision rate, and align the planning horizon to 3.0s for a fair comparison.
-We compute the $L2$ error between the planned trajectory and the human driving trajectory, and evaluate how often the SDV would collide with other agents on the road.
-Detailed comparison with previous methods is shown in Tab.
-Note that the Vanilla algorithm is penalized based on how much the trajectory deviates from the ground truth, thus it achieves the lowest L2 error but largest collision rates in all prediction horizons.
+For open-loop planning, our analysis prioritizes two pivotal evaluation metrics: the $L2$ error and the collision rate.
+The planning horizon is standardized at 3.0 seconds to ensure equitable comparison.
+We ascertain the $L2$ error by quantifying the deviation of the planned trajectory from the human driving trajectory and assess the frequency of potential collisions involving the ADV with other road agents.
+A comprehensive comparison with antecedent methodologies is delineated in Tab.
+2.
+It is noteworthy that the imitation strategy incurs penalties proportional to the trajectory’s divergence from the empirical ground truth, resulting in minimal $L2$ errors across all examined horizons.
 4.3 Closed-loop Experimental
-We first evaluate the necessity of different sampler on closed-loop simulation by ablation study.
-Then, find the best combination to evaluate in the real-world.
-The results of ablation study illustrated appendix.
-We extensively tested andn appendix.
-compare the above methods in the real world, in densely populated urban area, under the supervision of human safety drivers.
-The driving route used for our experiments is shown on appendix.
-During the 100+ mile public road testing, the model successfully performed a wide variety of challenging maneuvers including lane-following, merging, yielding to pedestrians or nudging around parked cars.
+Initially, we conducted an ablation study to assess the necessity of various samplers within a closed road environment.
+Subsequently, we identified the optimal sampler combination for further evaluation on open urban roadways.
+The outcomes of this ablation study are delineated in the appendix.
+Comprehensive real-world testing of these methodologies was conducted in densely populated urban settings, under the vigilant supervision of human safety drivers.
+The specific driving route utilized for our experiments is depicted in the appendix.
+Throughout the extensive public road testing, spanning over 30 kilometers, the model adeptly executed a diverse array of complex maneuvers, such as lane-following, merging, yielding to pedestrians.
+The results of these road tests are documented in Tab.
+3
+It is noteworthy that methodologies exhibiting superior imitation metrics do not invariably demonstrate enhanced performance in closed-loop evaluations.
+This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks.
+Nonetheless, the multi-modal imitation generator and GAN-based generator significantly enhance the system’s performance in both open-loop and closed-loop settings, primarily in terms of the planner’s intelligence and the continuity between consecutive frames.
+The efficiency and comfort metrics both show improvements with the integration of data-driven generators.
+Closed-loop results suggest that multimodality mitigates the brittleness inherent in imitation learning, while the lattice sampler further enhances the robustness of the entire system.
+To rigorously evaluate the efficacy of the proposed methodology as a substitute for off-the-shelf planners, we conducted a comparative analysis between the GAD and a conventional modularized, user-level, rule-based system, referred to as DPQP in Tab.
+3.
+Notwithstanding, a discernible performance discrepancy persists in terms of system’s robustness and comfortableness.
+We posit that increase training data size could ameliorate this disparity.
 5 Conclusions
+We introduce a sophisticated DL pipeline from prediction to motion planning in HD map-free setting, augmented with an enhanced safety assurance.
+Our investigations have yielded substantial enhancements over prevailing data-driven methodologies within the complex urban environments of Shanghai.
+This paradigm enables the safety deployment of learned planners in real-world scenarios, offering interpretability and leveraging their capacity to evolve with accumulating data, thereby managing scenarios of greater complexity than those addressed by traditional rule-based systems.
+Future work may integrate generative learning techniques such as Generative Adversarial Networks (GAN), Conditional Variational Autoencoders (CVAE), and diffusion models to further augment the data-driven capabilities in motion planning.
+Furthermore, evidence suggests that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners.
+It is recommended that subsequent research in motion planning incorporate the imitation metric to gauge the intelligence domain of planners and to develop additional metrics for closed-loop evaluation.
 6 Acknowledgements
@@
 Appendix A Appendix
-A.1 Scene Rasterization
+A.1 Perception Outputs
+At each timestamp, measured states for each tracked entity are provided in terms of positions, velocities, and accelerations.
+Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead.
+Stop lines are also essential when the ADV approaches the junction area.
+The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form.
+The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status, characterized as non-parametric.
+A.2 Scene Rasterization
 The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections.
+To capture the past motion of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers.
 Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect.
-To distinct SDV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history.
+To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history.
 The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness.
 we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route.
-Since we explore route from different source, we cast different routing form into one single uniformed *** All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
-A.2 Instance Vectorization
-A.3 Model Structure
-A.4 Route
-As indicated previously, the route plays a critical role in guiding SDVs along roadways.
+All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
+A.3 Instance Vectorization
+Parametric information such as surrounding vehicle positions, lane center lines and lane landmarks, are derived with the form of discretized points to depict the traffic scenes.
+Additionally, route information is obtained to facilitate the ADV in navigating to its intended destination.
+Given that the vehicle’s historical positions are temporally continuous, whereas the lane center lines, lane landmarks, and route information are spatially continuous, they are structured as vectors $\mathbf{v}$ that connect adjacent points.
+Specifically, the $i$th vector can be defined as
+$\mathbf{v}_{i}=[p_{i},p_{i+1},a_{i}]$ (17)
+where $p_{i}$ represents the $i$th point, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
+Similar to the process of scene rasterization, traffic lights are considered as static agents and can be vectorized using unchanged vectors.
+Each vector can be seen as a node in a graph.
+Subsequently, a sequence of nodes belonging to the same category will be combined to form an instance, which will then be interconnected within the graph.
+A.4 Model structure
+Our model architecture consists of two main branches.
+The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding.
+Subsequently, the output from the rasterization branch is fed into two identical-structured decoders, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch.
+The decoders generate cost volumes and prediction grid maps.
+Concurrently, the vectorization branch processes parametric instance information through a PointNet subgraph, which includes three edge convolutions.
+The output from this branch is then passed through a downstream Transformer encoder to obtain the instance encoding.
+Note that the rasterization branch has more layers than the vectorization branch as the input is much denser than the vectorized graph.
+Following the extraction of the scene embedding, features by oriented crop around the agents are obtained.
+These features are then processed through a sequence of max-pooling and a multi-layer perceptron (MLP).
+The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
+A.5 Route
+As indicated previously, the route plays a critical role in guiding ADVs along roadways.
 Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
-The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the SDV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
+The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
 However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes.
 Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions.
-Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the SDV to change lanes.
-Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the SDV lacking appropriate instructions to navigate intersections.
+Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the ADV to change lanes.
+Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections.
 To address this, the starting and ending positions for turning are recorded.
-As the SDV approaches intersections, human-driving trajectories are employed as the route to guide the SDV through the intersections.
-A.5 Curve Based Sampler
+As the ADV approaches intersections, human-driving trajectories are employed as the route to guide the ADV through the intersections.
+A.6 Curve Based Sampler
 In this appendix, we present the sampling procedure with more details.
+We follow [48] with little modifications for comfort concerns.
 Vehicle trajectories can be categorized into two primary types: straight and turning.
@@
 As the shapes of clothoid curves solely determined by the scaling factor a, we can generate a variety of clothoid trajectories by considering a suitable range for a, in combination with the initial curvature obtained from the bicycle model and the speed profiles.
-A.6 Lattice Sampler
+A.7 Lattice Sampler
 The lattice sampler is a reference line-based method used to acquire multiple candidate trajectories along the reference line, taking into account future movements of surrounding obstacles to avoid potential hazards.
@@
 If any trajectory does not meet the requirements, it will be removed from the sampling set.
-A.7 Retrieval-based Sampler
-The retrieval based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the SDV’s current state from a pre-established expert trajectory dataset.
+A.8 Retrieval-based Sampler
+The retrieval-based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the ADV’s current state from a pre-established expert trajectory dataset.
 To build the expert trajectory dataset, we utilized about 350 hours of driving trajectories from experienced human drivers.
@@
 To subdivide the original trajectories into different bins, we use the initial velocity v, acceleration a and curvature $\kappa$, with respective bin sizes of $1.0\frac{m}{s}$, $0.5\frac{m}{s^{2}}$, $0.01\frac{1}{m}$.
-In each bins, trajectories are clustered into 2000 sets and the closest trajectory to each cluster center is retained.
+In each bin, trajectories are clustered into 2000 sets, and the closest trajectory to each cluster center is retained.
 Finally, the trajectories from each subset will be aggregated to the comprehensive expert trajectory dataset.
@@
 […]
-$\textit{D}=\beta_{1}\cdot|v-v_{SDV}|+\beta_{2}\cdot|a-a_{SDV}|+\beta_{3}\cdot|\kappa-\kappa_{SDV}|$ (1)
+$\textit{D}=\beta_{1}\cdot|v-v_{ADV}|+\beta_{2}\cdot|a-a_{ADV}|+\beta_{3}\cdot|\kappa-\kappa_{ADV}|$ (1)
 In this paper, we assign different weights to the differences in velocity v, acceleration a and curvature $\kappa$, with $\beta_{1}=1.0$, $\beta_{2}=5.0$, $\beta_{3}=40.0$ and all nodes with $D<1.0$ will be added into output for the retrieval-based sampler.
-Figure 3: pipeline
-A.8 Hardware
+A.9 Hardware
 Our autonomous driving algorithms have undergone rigorous testing on the recently launched Jiyue 01 vehicle, whose detailed parameters can be found in the following link: https://www.jiyue-auto.com/robocar-01.
@@
 More NVIDIA-Orin parameters can be found in the following link: https://developer.nvidia.com/drive/agx.
-A.9 Task Perf
+A.10 Task Perf
+Figure 4: Tasks on board
+Task Time Cost(ms)
+Ave 90p Max
+Build Frame 0.77 1.13 5.47
+Build Obstacles 0.37 0.67 5.35
+Build Reference Line 0.52 0.82 10.61
+BackBone / Imitation planner 40.90 51.05 81.43
+Curve Based Sampler 5.12 10.75 15.64
+Retrival Based Sampler 8.31 12.16 19.22
+Lattice Sampler 9.83 15.15 40.51
+Cost Evaluator 0.58 1.39 9.95
+Safety Layer 0.32 0.53 8.12
+Over All 50.24 72.26 108.72
+Table 4: Task perf on board.
+The perception module is implemented on the master chip, whereas the GAD resides on the subsequent chip.
+We enumerate the time expenditures associated with each task in Tab.
+4.
+These tasks are orchestrated by a flow-based scheduler, facilitating concurrent execution as depicted in Fig.
+4.
+The deployment of the model via TensorRT is conducted without the utilization of acceleration techniques such as quantization or pruning.
 Figure 5: Test Drive Route
-A.10 Road Test
-Our road tests are conducted in the Jiading District of Shanghai, China, known for ”Shanghai Automobile City”.
+A.11 Road Test
+Our road tests are conducted in the Jiading District of Shanghai, China, known as ”Shanghai Automobile City”.
 The detailed testing route is illustrated in Figure 5, encompassing straight stretches, lane changes, intersections, and turning scenes involving various traffic participants.
-The route begins at No.1688 Yecheng Road, a road with two lanes in each direction, encounters frequent cars and electric bicycles.
+The route begins at No.1688 Yecheng Road, a road with two lanes in each direction that encounters frequent cars and electric bicycles.
 It passes through two intersections controlled by traffic lights before entering a three-lane structure and making a careful lane change to the right-most lane, where sometimes occupied by static vans.
@@
 Finally, the route circles back to Yecheng Road and returns to destination No.1688.
-A.11 Model structure
-Our model architecture consists of two main branches.
-The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding.
-Subsequently, the output from the rasterization branch is fed into two identical-structured decoders, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch.
-The decoders generate cost volumes and prediction grid maps.
-Concurrently, the vectorization branch processes parametric instance information through a PointNet subgraph, which includes three edge convolutions.
-The output from this branch is then passed through a downstream Transformer encoder to obtain the instance encoding.
-Following the extraction of the scene embedding, features by oriented crop around the agents are obtained.
-These features are then processed through a sequence of maxpooling and a multi-layer perceptron (MLP).
-The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
-A.12 Training details
-The Adam optimizer is used with an initial learning rate of $5xe^{-4}$ and a batch size of 32.
+The total length of the test route is 3.0 kilometers long.
+A.12 Dateset
+We collected the record files in 15 hours, which captured the perception output, the ADV trajectory, and route info.
+All the records are post-processed to obtain 200k clips in total.
+By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future.
+We partition the dataset into 16w for training and 4w for testing.
+We collected 15 hours of record files, containing messages such as perception outputs, ADV trajectories, traffic lights and route information.
+Following the collection, a thorough data cleaning process was implemented to analyze and identify the recorded data.
+Errors such as unconverged ADV localizations and perception frames of irregular frequencies were identified and excluded in order to enhance the quality and reliability of the dataset.
+The records were then categorized into different scenarios including cruising, lane-change, and turning scenarios.
+Subsequently, less-common but crucial scenarios, such as unprotected left turning scenarios, were amplified by data augmentation, in order to achieve a well-balanced dataset.
+While data augmentation resulted in higher loss during training since difficult scenarios were added to the dataset, it enhanced the model’s performance and generalization capability.
+Finally, all the records were post-processed to obtain 200k clips in total.
+By default, we take the 1.5 seconds of past contexts and plan for the future 3.0 seconds of contexts, corresponding to 15 frames in the past and 30 frames in the future.
+We partition the dataset into 160k for training and 40k for testing.
+A.13 Training details
+The Adam optimizer is used with an initial learning rate of $5e^{-4}$ and a batch size of 32.
 The scheduler of reduce on loss plateau decay is also implemented, with a reducing factor of 0.3 and a patience factor of 2, which means that if there is no improvement after 2 epochs, the learning rate will be reduced by 0.3 times the original value.
 The lower bound of the learning rate is set to $10^{-8}$ and the model parameters are randomly initialized.
-Experiments are conducted with 4 NVIDIA H800 GPUs with 60 epochs to complete.
-Add two stage training @weijian.sun
+Experiments are conducted with 4 NVIDIA A800 GPUs with 60 epochs to complete.
+In the first stage of our proposed two-stage training model, we regress a grid map by a cnn encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future.
+Specifically, cnn-encoder embeds the scene image into feature map with 12 convolution layers, and then 4 deconvolution layers are used to decode feature map to the output map.
+In the second stage of obstacle trajectory prediction, we reuse the CNN encoding layer from the first stage and initialize the encoder parameters with the last training step.
+As for obstacle trajectory prediction, we use an LSTM layer to replace the cnn decoder.
+In order to concentrate more on obstacles, we crop the encoding feature map centered agents, as the local scene feature of obstacles.
+The cropping direction aligns with the orientation of the obstacles.
+The local scene feature map will be embedding as a local scene vector through a Multi-Layer Perceptron layer and a max-pooling layer.
+Then, the local scene vector will be sent to the LSTM decoder, concatenated with agent’s history vectorized embedding vector.

### v3

@@
 Deep-learning-based techniques have been widely adopted for autonomous driving software stacks for mass production in recent years, focusing primarily on perception modules, with some work extending this method to prediction modules.
-However, the downstream planning and control modules are still designed with hefty handcrafted rules, dominated by optimization-based methods such as quadratic programming or model predictive control.
+However, the downstream planning and control modules are still designed with hefty hand-crafted rules, dominated by optimization-based methods such as quadratic programming or model predictive control.
 This results in a performance bottleneck for autonomous driving systems in that corner cases simply cannot be solved by enumerating hand-crafted rules.
 We present a deep-learning-based approach that brings prediction, decision, and planning modules together with the attempt to overcome the rule-based methods’ deficiency in real-world applications of autonomous driving, especially for urban scenes.
-The DNN model we proposed is solely trained with 10 hours of valid human driver data, and it supports all mass-production ADAS features available on the market to date.
-This method is deployed onto a Jiyue test car with no modification to its factory-ready sensor set and compute platform.
+The DNN model we proposed is solely trained with 10 hours of valid human driver data and supports all mass-production ADAS features available on the market to date.
+This method is deployed onto a Jiyue test car without modification to its factory-ready sensor set and compute platform.
 The feasibility, usability, and commercial potential are demonstrated in this article.
-[^1] Test records will be available at https://github.com/weijian-sun/GAD
 1 Introduction
@@
 However, this easily interpretable rule-based method is highly dependent on accurate prior knowledge, which severely restricts its scalability.
-High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information, greatly lessens the burden of perception and enormously facilitates planning tasks for autonomous driving.
+High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information.
 Unfortunately, HD maps are proven hard to scale up due to the complexity and costliness of map generation and maintenance, as well as failure to reflect changes in real-world road conditions timely, which is often constrained by resource-intensive collection processes.
@@
 New rules must be designed and re-tuned in accordance with different driving scenarios and new corner cases.
-This process is expensive and it scales poorly for the simple fact that the scenario space for the Autonomous Drive Vehicle(ADV) is infinite while rules can only be of limited number.
+This process is expensive and it scales poorly for the simple fact that the scenario space for the Autonomous Driving Vehicle(ADV) is infinite while rules can only be of a limited number.
 Figure 1: Comparison on the design of model-based planning.
@@
 (c)Proposed generative method, in which the data-driven generator gets feedback from the evaluator.
-End-to-end autonomous driving, which aims to consider sub-tasks in a full-stack manner jointly, has been proposed to resolve the above issues.
-Methods[22, 24, 25, 40, 20, 33, 5, 30] that direct learning from raw sensor data to planning trajectories or driving commands eliminate the need for hand-crafted heuristic rules and the prerequisite HD maps, avoiding infomation loss and better adapt to diverse scenarios.
-However, these approaches lack interpretability which plays a key role in understanding and tracing how the system envolves, thus will be difficult for the developer to optimize.
-Interpretability is of key importance in a safety-critical system, particularly if a bad event were to happen in mass production.
-Moreover, the absence of modularized structure and prior knowledge makes these methods very brittle when deployed in real-world closed-loop.
+End-to-end autonomous driving, which aims to consider sub-tasks in a full-stack manner jointly, has been advocated to address the aforementioned challenges.
+Methods[23, 25, 26, 43, 21, 35, 5, 32] that facilitate direct learning from sensor data to trajectories or driving commands obviate the necessity for manually devised heuristic rules and the dependency on HD maps.
+However, these methodologies are deficient in interpretability, which is crucial for elucidating and monitoring the system’s evolution, consequently complicating optimization efforts by developers.
+Moreover, the absence of modularized structure and prior knowledge makes these methods exceedingly brittle when operationalized in real-world closed-loop.
 To overcome the aforementioned problems, we explore an alternative method that runs entirely on the edge and purely relies on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [13] based on perception results.
 The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first, primarily due to considerations of scalability and mass production feasibility.
-We believe this is effective for two primary reasons.
-Firstly, landmarks provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized.
-Secondly, for turning scenes, we gather routing information from historical commuting trajectories and avoid collisions with the help of real-time occupancy.
 However, the proposed methodology is not limited to commuting scenarios; it supports a versatile selection of route formats.
 We also validated its non-commuting variant in the experiments.
-To break the bottleneck brought by rule-based planning but maintain the system’s explainability and stability, we propose a general framework to handle the prediction-planning tasks in a data-driven and integrated manner.
+To break the bottleneck brought by rule-based planning, we propose an interpretable and tunable framework to handle the prediction-planning tasks in a data-driven and integrated manner.
 Our framework consists of a trajectory generator and evaluator.
-The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric info in grid form to validate and score candidates [^2] We note that by parametric form we mean the structured numerical representations of certain perception outputs, such as the observed vehicle states.
+The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric information in grid form to validate and score candidates [^1] We note that by parametric form we mean the structured numerical representations of certain perception outputs, such as the observed vehicle states.
 Conversely, the non-parametric form refers to unstructured representations that cannot be described by parameters, such as occupancy..
 We validate our framework in a closed loop by deploying and testing it in real-world user-level autonomous driving vehicles.
-We elucidate the rationale behind our architectural that segregates perception from the comprehensive end-to-end stack: Firstly, for mass production friendly concern, the modularized perception provides explaninablity and tunablitiy, Secondly, information articulated in parametric form facilitates reasoning at the agent level, not solely for application within the safety layer and the analysis of explainable cases, but also to enhance our trajectory generation processes.
-Thirdly, parameterization encompasses post-tracking and fusion for the results of single or multi-modal detection on raw sensor data which remain formidable unresolved issues, particularly when considering the substitution of these processes with a model to attain comparable results.
-We leave perception involved version to follow up works when those issue resolved in the future.
-The real-world closed-loop validation needs to be highlighted as well: Firstly, a on road planer will adjust to the target in the loop to keep safe and comfortable since real-world planning needs handle the deviations brought by the external disturbances and controller’s modeling errors, which is hard to simuate and omitted by open-loop planner as well.
-Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller.
-Since unexpected and unreasonable jerks and curvatures can only be tested when the controller is in the loop with the planner.
-The above limitations make the results in open-loop or simulator less convincing and leaves the effectiveness of those methods (like UniAD[22], VAD[24], mile[20], ST-P3[21]) unknown for onboard road tests and mass production.
-Our key contributions are summarized as follows:
-  1. 1.
-    We propose an industry-level paradigm for data-driven prediction-planning with an HD map-free setting.
-Our model leverages non-parametric information with better tolerance to upstream perception and handling prediction and planning tasks in one backbone.
-  2. 2.
-    To the best of our knowledge, we are the first evaluation of such system in a complex, real-world, urban road with a factory-ready sensor set and compute platform.
+Our key contributions are summarized as follows: (1) We introduce a scalable, tuneable, industry-grade framework for data-driven prediction-planning in HD map-free setting.
+Our model offers an alternative to the e2e model and demonstrates performance on par with off-the-shelf, optimization-based planner.
+(2) To the best of our knowledge, we are the first evaluation of such system in a complex, real-world, urban road with factory ready sensor set and compute platform [^2] Road test video is available at https://youtu.be/DYPB5lCRvn0.
 We demonstrate the imperative of closed-loop evaluation compared to imitation metric adopted by extensive offline studies.
-  3. 3.
-    We extend the state-of-the-art data-driven approach and propose a method that goes beyond simple imitation or trajectory sampling and instead reasons about the potential in generative learning by combining max marge planning with multi-modal imitation.
-As depicted in Fig.
+(3) We extend the state-of-the-art data-driven approach by proposing a method that goes beyond simple imitation or trajectory sampling.
+Instead, our method reasons about the potential in generative learning by combining max margin planning with multi-modal imitation, as depicted in Fig.
 1
@@
 2.1 End to End Learning&Validation
-End-to-end autonomous driving has become a significant research focus, with numerous recent studies.
-These methods, using neural networks, produce planning trajectories or control commands, with many[40, 47, 20] focusing on closed-loop simulator experiments [10, 27].
-However, a persistent domain discrepancy exists between the simulated environment and the actual world, especially in terms of sensor data fidelity and the dynamic states of agents.
-The accurate modeling of agent behaviors and the simulation of real-world data[14] within closed-loop environments continue to pose formidable challenges that have yet to be resolved.
+End-to-end autonomous driving has become a significant research focus.
+These methods, using neural networks, produce planning trajectories or control commands, with many[43, 50, 21] validations on closed-loop simulator[11, 29].
+However, a persistent domain discrepancy exists between the simulated environment and the actual world, the accurate modeling of agent behaviors and the simulation of real-world data[15] within closed-loop environments continue to pose formidable challenges that have yet to be resolved.
 Recently, open-loop experiments with real-world data have attracted more attention[25, 23, 22, 51].
 Methods like UniAD[23] claim their effectiveness in improving final planning performance by involving learning intermediate tasks.
-However, it relies on an additional optimal control module to obtain a feasible trajectory and prevent collisions, which violates the initial purpose of breaking optimization-based deficiencies in motion planning task.
-Nevertheless, the effectiveness of open-loop demonstration has been challenged by [28] in terms of imbalanced data distribution and metric limitation.
-We address these contentious issues by rigorously validating our approach through real-world closed-loop road test, wherein the proposed methodology eschews the dependence on conventional optimization algorithms to derive trajectories that adhere to specified constraints.
+Nevertheless, it necessitates an auxiliary optimization-based module to generate a viable trajectory and avert collisions, thereby contravening the foundational objective of circumventing deficiencies inherent in optimization-based approaches.
+Moreover, the effectiveness of open-loop demonstration has been challenged by [30] in terms of imbalanced data distribution and metric limitation.
+Those unresolved limitations make the results in open-loop or closed-loop simulator less convincing and leaves the effectiveness of methods (like UniAD[23], VAD[25], mile[21], ST-P3[22]) unknown for onboard road tests and mass production.
+We address these contentious issues by rigorously validating our approach through real-world closed-loop road test, wherein the proposed methodology eschews the dependence on conventional optimization-based algorithms.
 2.2 Data Driven Prediction
-Prior research has extensively explored actor-based prediction methods for future driving dynamics.
-These methods either produce specific trajectories [4, 2, 32, 51] or use generative models to outline probable outcomes [36, 41, 17, 26].
-In order to encode scene context, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8].
+Prior researches have extensively explored actor-based prediction methods for future driving dynamics[4, 2, 34, 54, 38, 44, 18, 28].
+In order to encode scene contexts, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8].
 More recent methods have started to leverage the topologic graph obtained from HD maps in order to better represent lane connectivity.
 VectorNet [14] encodes map and agent as polylines with a global interaction graph.
-LaneGCN [29] treats actor past and the lane graph separately and then fuses them with a series of attention layers.
+LaneGCN [31] treats actor histories and the lane graph separately and then fuses them with a series of attention layers.
 Temporal occupancy maps [24, 39, 20], on the other hand, provide another option to forecast grid or heat map at the scene level.
 We follow the philosophy of [12, 19, 41, 45, 49] in modeling scene-level representations but propose a multi-stage pipeline to perform forecasting in both formats.
-Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are mainly adopted in the safety layer and to boost some of our samplers.
+Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are primarily used to enhance the safety layer and to boost some of our samplers.
 2.3 Data Driven Planning
-For imitation based methods, modeling strategies from trajectory prediction are adopted, which build an imitation decoder trained under the supervision of human demonstration.
-Apart from the aformention end to end works, SafetyNet[44] developed a machine learning planner and tested it in downtown San Francisco, but the model can hardly avoid collision without the help of a fallback layer.
+Apart from the aformention end to end works, imitation-based methods have been widely explored in the planning task.
+SafetyNet[47] developed a machine learning planner and tested it in downtown San Francisco, but failed to avoid collision without the help of a fallback layer.
 VAD[25] modeled the driving scene as a fully vectorized representation.
 PlanT[37] proposed a transformer-based backbone in actor-level representation and validated in simulation.
-However, the monotonous single trajectory imitation suffers from robustness and lacks interpretability[6], thus difficult to tune by case and hard to scale up.
-More recently, explicit methods with great interpretability have been shown to adapt better to more challenging environments.
-Such approaches usually build a cost volume with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost.
-The cost volume is organized in a non-parametric form learned directly from the network[48, 38, 3, 49, 7].
-Beginning with pioneering work on NMP [48], DSDNet[49] and P3[38] made further developments that combined hand-made and learning-based costs to obtain an integrated cost volume.
-In the follow-up work, MP3 [3] introduced a mapless approach, while ST-P3[21] explored vision-based inputs.
-We also adopt this combination to choose the best trajectory but extend the system’s robustness by boosting the generating algorithm with modularized perception output.
-In contrast to all methods above, which solely rely on a rule-based sampler, our model leverages multi-modal imitation learning to build a diversified hybrid trajectory set and stays safe by being evaluated in an interpretable cost volume as a summary of scene occupancy and agent motion.
-Figure 2: Model Structure of GAD.
-The model initiates with two streams for rasterization and vectorization respectively to encode scene and instance-level features.
-The rasterization branch utilizes convolution layers to derive the scene embedding, which is then decoded into cost volumes for scoring candidate trajectories from multiple samplers and prediction grid maps for agent prediction.
-Meanwhile the scene embedding are cropped and incorporated with the instance encoding from vectorization branch to decode ego trajectory and agents’ future movements.
+However, the monotonous trajectory regression suffers from robustness and lacks interpretability[6], thus difficult to tune by case and hard to scale up.
+More recently cost-volume-based methods have improved adaptability and interpretability in challenging environments.
+They identify optimal trajectories by evaluating candidates within a network-learned cost volume[51, 41, 3, 52, 7], while the candidate trajectories are generated through rule-based samplers.
+Beginning with pioneering work on NMP [51], DSDNet[52] and P3[41] made further developments that combined hand-made and learning-based costs.
+In the follow-up works, MP3 [3] introduced a mapless approach, while ST-P3[22] explored vision-based inputs.
+We also adopt this combination to choose the best trajectory, but in contrast to all methods above, which solely rely on a rule-based sampler, we explore the potential of multi-modal imitation and generative learning in motion planning task.
+Figure 2: The Model Structure of GAD initiates with two streams, rasterization and vectorization, to encode scene and instance-level features; the rasterization branch uses convolution layers to derive the scene embedding, which is decoded into cost volumes and prediction grid maps, while the scene embedding is cropped and combined with the instance encoding from the vectorization branch to form $f_{agent}$ which decoded into ego trajectory and agents’ future movements.
 3 Methods
@@
 In our work, the perception module follows a typical mainstream design that provides tracked entities, landmarks, and occupancy voxels (also known as volumn pixels) by incorporating images from a suite of independent cameras with different field of views (FoVs) to formulate a surround view at 360 degrees.
-The perception module has a deep neural net that directly stitches features in bird’s-eye view (BEV) form, and transformers are used to capture both spatial and temporal contexts.
-Occupancy is also an integral part of outputs by this perception module, providing a useful scene representation of a three-dimensional world that offers free space for SDV to navigate through.
-More details are provided in the appendix.
-To encode upstream output for different usages, we design representations for scene and instance levels [^3] We note that by scene we mean the overall environment context for the current frame, by instance we mean the elements in the scene such as ego, agent, landmark, route, etc., which are applied to trajectory evaluation and generation, respectively.
+More details are provided in Sec.
+A.1.
+To encode the upstream output for diverse applications, we have devised representations at both the scene and instance levels [^3] Here, by scene, we refer to the comprehensive environmental context pertinent to the current frame, and by instance, we denote the constituents within the scene such as the ego vehicle, other agents, landmarks, routes, etc., which are systematically utilized in the evaluation and generation of trajectories.
 Scene Representation
-We adopt rasterization proposed by[9] for scene representation.
-Scene rasterization with convolutional forwarding can better capture scene-level cues compared to instance vectorization, avoiding information loss to the maximum extent.
-Owing to the convolutional network’s property of exploiting spatially local correlation, undefined or irregular elements like construction areas, traffic cones, and roadblocks can be incorporated with a unified design regardless of how they are organized and whether they have regularized shapes.
-Additional scene information depicted by free space or occupancy, lacking parametric description, can hardly be involved with vectorized representation and, once omitted, can be fatal to safety-critical systems.
-To serve as a qualified trajectory evaluator, BEV rasterization involves environment cues from online perception outputs as much as possible to capture those “devils in detail”.
-We follow the interpretable design as that of in [3], but rasterize different scene elements(e.g.
-agent, lanes, occupancy, route) into different binary channels to enable reasoning about the distinct element separately.
-The details can be found in our appendix.
+We employ the rasterization methodology as proposed by [10] for the representation of scenes.
+Rasterization, when coupled with convolutional forwarding, more effectively captures scene-level cues than instance vectorization.
+Due to the inherent ability of convolutional networks to exploit spatial correlations, it is feasible to integrate undefined or irregular elements such as construction zones, traffic cones, and roadblocks into a cohesive design, irrespective of their arrangement or geometric regularity.
+Furthermore, scene information characterized by free space or occupancy, which lacks a parametric framework, is challenging to incorporate within a vectorized representation and, if excluded, poses significant risks to safety-critical systems.
+BEV rasterization assimilates environmental cues from online perception outputs to the fullest extent to capture those ’devils in detail’.
+Comprehensive details are provided in Sec.
+A.2.
 Instance Representation
-The instance representation is employed to address the shortcomings of scene encoding, which are afflicted by the truth that, rasterization is prone to distorting highly organized and structured information, rendering it disorganized.
-Since this procedure is irreversible, making it challenging to retrieve instance encoding for lane center lines, road boundaries, and other elements.
-The vectorized representation combined with the transformer has been proven to be superior to CNN forwarding on BEV raster in terms of accuracy and efficiency[13] for the motion prediction task.
-To take advantage of modeling in such form, we adopt vectorization[13] to encode the perception outputs into vector sets.
-Instances with parametric description are encoded in the vector form, including route points, ego and agent historical, lane center lines, and road boundaries, while non-parametric outputs are ignored, like occupancy, construction area, roadblocks, and other irregular elements.
-Each element includes position and type features, with the position described relative to the ADV pose.
-The relationships between separated elements are organized as a graph.
+The instance representation is utilized to mitigate the deficiencies inherent in scene encoding, which suffers from the fact that rasterization tends to distort highly organized and structured data into non-parametric forms.
+Given that this process is irreversible, it complicates the extraction of instance encoding for elements such as lane centerlines, road boundaries, among scene feature map.
+On the other hand, vectorized representation combined with the transformer has been proven to be superior to CNN forwarding on BEV raster for instance reasoning[14].
+To leverage the modeling capabilities afforded by this form, instances with parametric description are encoded in the vector form, including route points, agent histories, lane center lines, and road boundaries, while non-parametric outputs are ignored, such as occupancy, construction area, roadblocks, and irregular elements.
+More details can be found in Sec.
+A.3
 3.2 Model Structure
@@
 One steam processes the rasterized scene while the other deals with the vectorized graph.
-To elaborate, the rasterization stream employs a convolutional encoder of multiple resolutions to derive the scene embedding, responsible for aggregating geometric information from the scene context to extract global appearance and motion cues.
-The scene embedding is then passed through two separate deconvolution decoders to generate the cost volumes and the predicted grid maps respectively.
+To elaborate, the rasterization stream employs a convolutional encoder of multiple resolutions to derive the scene embedding, which is then passed through two separate deconvolution decoders to generate the cost volumes and the predicted grid maps respectively.
 Meanwhile, in the other stream, inspired by [14], the instance-level encoding is built on a hierarchical graph network-based architecture.
@@
 Then an imitation decoder is utilized to produce the trajectories for agents and the ego.
-The details for the encoders and decoders are illustrated in the appendix.
+The details structure for the encoders and decoders are illustrated in Sec.
+A.5.
 3.3 Route
-When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the ADV should follow.
-However, in HD map-free settings, this is not possible, especially for turning scenarios in intersections where no leading landmark is available.
-In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues.
-The lane center lines constructed by landmarks are mainly used in cruise and lane change scenarios, while in turning scenarios, the starting and ending positions are recorded in advance, once vehicles approach the intersection, the human commuting trajectories will serve as route.
-More details can be found in the appendix.
+When HD maps are available, the route is typically a sequence of lanes for the ADV to follow.
+In settings without HD maps, this approach is not feasible.
+In this work, route information is derived from visual landmarks and human commuting trajectories for different scenarios to substitute those absent cues.
+This approach is deemed effective for two primary reasons.
+Firstly, landmarks (lane center, lane divider and road boundary) provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized.
+Secondly, for turning scenes(left turn, right turn, u-turn), we gather routing information from historical commuting trajectories, supplemented by real-time occupancy data to mitigate collision risks.
+The implementation details can be found in Sec.
+A.6.
 3.4 Prediction
-We follow a multi-stage design for scene and agent prediction.
-In the first stage, the pipeline aims to predict the scene level grid map, also known as the heat map or occupancy.
-We regress a grid map $G_{ij}^{t}=dec(f_{ras})$ with a resolution matches that of the rasterization input by decoding the middle layer features.
-Figure 3: Multi-stage Prediction
-The decoder, as mentioned in Sec.
-3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box.
-This grid is trained with pixel-wise focal loss inspired from [15]:
-$\begin{array}[]{r}L_{occ\_pred}=-\sum_{ij}[\tilde{g_{ij}}(1-\hat{g_{ij}})^{2}log(\hat{g_{ij}})+\\ (1-\tilde{g_{ij}})\hat{g_{ij}}^{2}log(1-\hat{g_{ij}})]\end{array}$ (1)
-Here, $\tilde{g_{ij}}$ represents the class label for the grid, where $\tilde{g_{ij}}=1$ indicates that it is occupied.
-In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid.
-For the agent-level prediction, we follow the interactive design.
-Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego.
-In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior.
-Those landmarks indicate areas where interaction is most likely to happen.
-The agents’ trajectories will be decoded conditioned on that lane prior:
+We employ a multi-stage design for scene and agent prediction as depicted in Fig.
+3, where the first stage, involves predicting the scene-level grid map (also known as the heat map or occupancy) by regressing $G_{rid}=dec(f_{ras})$ with a resolution that matches the rasterization input through decoding of the scene embeding.
+Temporal resolution $t$ is 0.1 second from 0 to 3 seconds.
+The grid prediction decoder, as mentioned in Sec.
+3.2, shares a similar structure with the cost volume decoder, yet it is supervised from the rasterization of the agents’ future ground truth box.
+This grid is trained with pixel-wise focal loss inspired from [16], for brief we replace $G_{rid}$ with $g$:
+$\begin{array}[]{r}L_{pred\_grid}=-\sum_{ijt}[\tilde{g_{ijt}}(1-\hat{g_{ijt}})^{2}log(\hat{g_{ijt}})+\\ (1-\tilde{g_{ijt}})\hat{g_{ijt}}^{2}log(1-\hat{g_{ijt}})]\end{array}$ (1)
+Here, $\tilde{g_{ijt}}$ represents the class label for the grid, where $\tilde{g_{ijt}}=1$ indicates that it is occupied in timestamp t.
+In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid by following an interactive design.
+Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, it captures only nearby agents, assuming interactions with the ego.
+In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas $P_{lane}=\sum_{t}\sum_{i,j\in(lane)}g_{ij}^{t}$, filtering out higher probability candidates as the target prior.
+These landmarks indicate likely interaction areas, and agents’ trajectories are decoded based on this lane prior:
 $Traj=dec(f_{agent}|f_{lane})$ (2)
 Here, $f_{agent}$ is the instance encoding as illustrated in Sec.
-3.2, $f_{lane}$ is the corresponding lane encoding obtained in the same way.
+3.2, $f_{lane}$ is the corresponding lane encoding obtained in the vectorized stream.
 For the training phase, the ground truth $\tilde{lane}$ corresponding to the agent’s future trajectory will be labeled and incorporated as a condition prior:
@@
 The training loss is measured by the $L2$ distance between the predicted and ground truth trajectory.
+Figure 3: Multi-stage Prediction
 3.5 Trajectory Evaluation
-The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories.
-We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost volumes for scoring candidate trajectories.
-We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map.
-Our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster.
-The cost of a trajectory is defined as the sum of costs at its space-time way-points.
-The best candidate trajectory is the one with the lowest cost according to the learned cost:
+The previous section introduced a multi-stage prediction method that produces grid maps and interactive trajectories.
+We now demonstrate the grid can be seamlessly integrated with perception occupancy into an end-to-end motion planner utilizing space-time cost volumes to evaluate candidate trajectories.
+Following the framework of [51] for max-margin planning, we refine it to include scene occupancy and anticipated grid maps, and our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form for future 30 consecutive frame from BEV raster.
+The trajectory cost is calculated as the sum of costs at its space-time waypoints, with the best trajectory being the one that minimizes this cost:
 $s=arg\min_{\hat{s}\in S}\sum_{t}[C(\hat{s})+\alpha O_{cc}(\hat{s})+\beta G_{rid}(\hat{s})]$ (4)
-where $S$ represents the set of viable future trajectories, $O_{cc}$ represents perception occupancy, $G_{rid}$ is the predicted grid map, $\alpha$ and $\beta$ are the weights respectively.
-Since selecting the minimum-cost trajectory within a discrete set is not differentiable, we use the max-margin loss to penalize trajectories that have small cost and are different from the human driving trajectory or are unsafe.
-Let $\hat{s}$ and $\tilde{s}$ be the input and human trajectory respectively for a given example.
-We utilize the max-margin loss to encourage the human driving trajectory to have a smaller cost than other trajectories.
+where $S$ represents the set of planed trajectories, $O_{cc}$ represents perception occupancy, $G_{rid}$ is the predicted grid map, $\alpha$ and $\beta$ are hyperparameters.
+Since selecting the minimum-cost trajectory within a discrete set is not differentiable, we use the max-margin loss to penalize trajectories that either deviate from the human driving trajectory or are unsafe, and to encourage the human driving trajectory $\tilde{s}$ to have a smaller cost than other trajectories $\hat{s}$ for a given example:
 $L_{cost\_volume}=\max_{\hat{s}}[C(\tilde{s})-C(\hat{s})+Diff(\tilde{s},\hat{s})]$ (5)
-The $Diff$ value here serves as compensation to quantitatively measure the difference between the human demonstrated and our candidates.
-It has to tell how goodness of the human driving or how badness of our samplers, quantified by planning targets, avoiding static collisions, and managing dynamic interactions, etc.
-So that the evaluator can assign comparable costs to the trajectories in our candidate set.
+The $Diff$ metric functions as a compensatory quantifier, meticulously gauging the disparity between human demonstrations and algorithmic candidates.
+It is designed to assess the efficacy of human driving and the deficiencies of candidates, as delineated by objectives such as planning target, evasion of static obstacles, and regulation of dynamic interactions.
+This enables the evaluator to attribute commensurate costs to the given trajectories.
 In particularly:
-$Diff(\tilde{s},\hat{s})=L2(\tilde{s},\hat{s})+O(\hat{s})+G(\hat{s})$ (6)
+$Diff(\tilde{s},\hat{s})=L2(\tilde{s},\hat{s})+O(\hat{s})+G_{rid}(\hat{s})$ (6)
 where $O$ is the occupancy cost, $G$ is the rest of the prediction costs as defined in Sec.
@@
 Note, however, that this optimization with physical constraints is NP-hard.
-On one hand, we evaluate and extend previous works on trajectory sampling[48, 3, 38, 21] with a more sophisticated design by incorporating prior knowledge and structured outputs from perception and prediction.
-Furthermore, we leverage a GAN-based pipeline during the generator’s training to solve the optimization problem.
+On one hand, we evaluate and extend previous works on trajectory sampling[51, 3, 41, 22] by incorporating prior knowledge and structured outputs from perception and prediction.
+Furthermore, we leverage a GAN-based pipeline during the imitation generator’s training to solve the optimization problem.
 In this section, we describe our sampling and generation algorithm.
-The sampling and generating algorithms can be mainly categorized into reference line-based and reference line-free.
-Methods incorporating reference information can generate long-term consistent outputs in the loops.
-Since this information rarely changes over time, bridging consecutive frames promotes cross-frame stability.
-However, scenarios still need to be handled where reference-related information can hardly be obtained.
-To enhance the scalability of our system, we follow [48, 3] to build curve and retrieval-based sampler in reference-line free and illustrate them in the appendix.
-The proposed lattice sampler, imitation planner, and GAN-based planner will be explained in this section.
-They incorperate reference info constructed by visual landmarks or commuting trajectories.
+The sampling and generating algorithms are categorized into reference line-based and reference line-free types, where methods that incorporate reference information provide long-term consistent outputs and enhance cross-frame stability as this information remains largely unchanged over time.
+However, challenges arise in scenarios where reference-related information is scarce.
+To improve our system’s scalability, we follow [51, 3] to develop curve and retrieval-based samplers in a reference-line free context, detailed in Sec.
+A.8 and Sec.
+A.9.
+Additionally, this section will discuss the proposed lattice sampler, imitation planner, and GAN-based planner, which utilize reference information constructed from visual landmarks or commuting trajectories.
 Lattice Sampler
@@
 When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process.
-The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in longitude.
+The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in the longitudinal direction.
 In our paradigm, they are both available as depicted by Sec.
@@
 In particular, we follow the trajectory parameterization and sampling philosophy proposed in [48, 42].
-Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude.
-We illustrate the detail in the appendix.
+The sampled trajectories include maneuvers such as lane keeping, lane changing, nudging, side passing in the lateral direction, as well as following, accelerating, yielding, and stopping in the longitudinal direction.
+More details can be found in Sec.
+A.10.
 Imitation Planner
 For the imitation planner, previous approaches choose to mimic a human driver with one single trajectory[23, 25].
-But such designs suffer from model collapse in both longitude and lateral.
+But such designs suffer from model collapse in both longitudinal and lateral directions.
 Owing to our generator-evaluator paradigm, we can generate a bunch of diverse candidates and pass them to the downstream evaluation task.
-We follow multiple-trajectory[8] design in trajectory prediction to estimate the conditional distribution $P(S|f_{plan})$ in motion planning task.
+We follow multiple-trajectory[8] design to estimate the conditional distribution $P(S|f_{plan})$ and multi-modality in motion planning task:
+$P(S|f_{plan})=\sum_{M}P(S|m,f_{plan})P(m|f_{plan})$ (7)
 Where $S=[\hat{s_{1}},...\hat{s_{m}}]$ is the trajectories with $m$ different maneuvers such as lane following, lane changing, agent following, agent overtaking.
-$f_{plan}$ stands for feature prepared for planning task.
-||||| | stands for concatenation.
-$f_{plan}=[f_{scene}||f_{route}||f_{ADV}]$ (7)
-$P(S|f_{plan})=\sum_{M}P(S|m,f_{plan})P(m|f_{plan})$ (8)
-Our imitation generator decodes a multitude of $M$ possible trajectories for ego move in local coordinates.
+$f_{plan}$ is constructed by feature prepared for planning task in Eq.
+8, where (||||| |) stands for concatenation.
+$f_{plan}=[f_{scene}||f_{route}||f_{ADV}||f_{target\_speed}]$ (8)
+The imitation generator decodes $M$ possible trajectories for ego move in local coordinates. During training, we then identify modal closest to the ground-truth trajectory according to $L2$ distance.
 ${[\hat{s_{1}},...\hat{s_{m}}]}=G(f_{plan})$ (9)
-During the training phase, we then identify mode $m^{\ast}$ that is closest to the ground-truth trajectory according to distance function,
 $L_{imit\_plan}=arg\min_{i\in{1,...m}}dist(\tilde{s},\hat{s_{i}})$ (10)
-We select $L2$ dist for the distance function.
-GAN Based Planner
-The above imitation generator is trained separately from our evaluator by different headers.
-With the attempt to generate trajectory with minimal cost, we propose a GAN base planner with modification to the training paradigm that bridges the isolated optimization targets, so that Eq.
-4 can be solved by learning.
-For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize the following function while the discriminator tries to maximize it.
+GAN-Based Planner
+The aforementioned imitation generator is trained independently from our evaluator under distinct supervisory decoder.
+In pursuit of generating trajectories utilizing the acquired cost volume and addressing Eq.
+4 through a learning-based approach, we propose a GAN-based planner with a conditional generative adversarial training framework.
+This framework is designed to bridges the isolated optimization targets between generator and evaluator, thus facilitating the learning of a stochastic generative model beyond trajectory imitation, adept at encapsulating the multi-modal diversity inherent in planning trajectories.
+Regarding the standard adversarial loss[17], commonly referred to as the min-max loss, the generator endeavors to minimize the subsequent function, while the discriminator seeks to maximize it.
 $L_{gan}=\min_{G}\max_{D}(E_{x}[log(D(x))]+E_{z}[log(1-D(G(z)))])$ (11)
 $log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly.
-In our paradigm, we replace the discriminator with an evaluator.
-But unlike the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator outputs a score to measure the discrepancy between candidates quantitatively.
-So the optimization target for evaluator Eq.
-5 still holds, but the learning target of the imitation planner becomes:
-$\ L_{generator}=E(G(m))+Diff(\tilde{s},G(m))$ (12)
+In our proposed framework, the conventional discriminator is supplanted by the evaluator as delineated in Sec.
+3.5.
+Contrary to the traditional discriminator in GANs, which categorizes instances with qualitative labels denoting veracity, our evaluator quantifies the divergence among candidates through the learned cost volume.
+Thereby, the optimization target for evaluator Eq.
+5 still holds, but the learning target of the generator changed.
+The complete formula derivation is provied in Sec.
+A.11.
+$\ L_{generator}=E(G(m))+L_{imit\_plan}(\tilde{s},G(m))$ (12)
 $L_{evaluator}=\max_{\hat{s}}[E(\tilde{s})-E(\hat{s})+Diff(\tilde{s},\hat{s})]$ (13)
 with:
-$E(s)=C(s)=\sum_{(x_{t},y_{t})\in{s}}C(x_{t},y_{t})$ (14)
-$\ C_{xy}^{t}=dec(f_{ras})$ (15)
-$\ G(m)=\hat{S}=[\hat{s_{1}},...\hat{s_{m}}]$ (16)
-Here, m is a hidden variable representing the manual in planning.
+$E(s)=\sum C(s)+\alpha O_{cc}(s)+\beta G_{rid}(s)$ (14)
+$[\hat{s_{1}},...\hat{s_{m}}]=\hat{S}=G(m);m\sim\mathcal{N}(\mu,\,\sigma^{2})\,.$ (15)
+The architecture is exactly the same as presented in previous subsections, except that in the generative model, the hidden variable $m$, which indidates driving manual in deterministic model is sampled from a Gaussian distribution.
+The final encoding for ADV is concatenated with $m$ to decode into its trajectory.
 However, the trajectory evaluation Eq.
@@
 3.7 Safety layer
-For on-road autonomous driving motion planning, a strategy for safety coverage is necessary.
+For on-road autonomous driving, a strategy for safety coverage is necessary.
 The perception occupancy and trajectory prediction results are used to validate the trajectory candidates with respect to safety.
@@
 4 Experiments
-We develop various baseline variants of our models under both open-loop and closed-loop configurations for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies.
+We develop various baseline variants of our models under both open-loop and closed-loop experiments for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies.
 We gather in-house datasets for both training and open-loop evaluations, encompassing an extensive array of driving scenarios within densely populated urban settings.
@@
 4.1 Prediction
+Method Mean L2 / hit-rate $<$ 1m Avg.
+1s 2s 3s VectorNet[14] 0.75/0.80 1.72/0.40 2.85/0.22 1.34 GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35 TNT[53] 0.65/0.78 1.01/0.57 1.72/0.34 0.90 GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
+Table 1: Prediction Results in uni-modal and multi-modal.
 For the prediction task, prior state-of-the-art methodologies are appraised within HD map-free context, wherein solely the observable landmarks are furnished to the predictors.
@@
 In general, our multi-modal design with interactive prior outperformed our reimplemented TNT[53] with 16 candidates, while the uni-modal version shared similar results with the current state-of-the-art.
-Method Mean L2 / hit-rate $<$ 1m Avg.
-1s 2s 3s
-VectorNet[13] 0.75/0.80 1.72/0.40 2.85/0.22 1.34
-GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35
-TNT[50] 0.65/0.78 1.01/0.57 1.72/0.34 0.90
-GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
-Table 1: Prediction Results in uni-modal and multi-modal.
-Method trajectory sources L2(m) Collision(%)
-curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s
-UniAD[22] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78
-VAD[24] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69
-NMP[48] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05
-MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97
-P3[38] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04
-GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52
-GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48
-GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07
-GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
-Table 2: Open loop planning results of imitation based method, sampling based method and GAD.
-Method comfort event in 30Km efficiency MPI
-Lat. accel. Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute
-UniAD[22] 0.351 3.12 3.68 - - - - 15.20 0.31
-ST-P3[21] 0.857 4.54 8.53 4 3 14 3 8.62 1.36
-GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72
-GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75
-GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75
-GAD-GAN 0.511 1.15 3.52 3 1 2 1 5.51 4.28
-DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
-Table 3: Road Test Outcomes: Data were systematically collected during a 30 km road test, encompassing the computation of average time per loop.
-Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded.
-The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
 4.2 Open-loop Planning
@@
 A comprehensive comparison with antecedent methodologies is delineated in Tab.
+3.
+It is noteworthy that the imitation strategy incurs penalties proportional to the trajectory’s divergence from the empirical ground truth, resulting in minimal $L2$ errors across all examined horizons.
+Method hist tar v L2 passing rate
+ras vec m/s Ave straight turn
+GAD-multi_1 $\surd$ $\surd$ x 0.56 0.00 0.00
+GAD-multi_2 $\surd$ $\surd$ $\surd$ 0.63 0.80 0.75
+GAD-multi_3 $\surd$ x $\surd$ 0.70 0.85 0.70
+GAD-multi_4 x x $\surd$ 0.34 0.65 0.00
+Table 2: Ablation study on ego feature for imitation planner.
+Method trajectory sources L2(m) Collision(%) curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s UniAD[23] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78 VAD[25] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69 NMP[51] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05 MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97 P3[41] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04 GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52 GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48 GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07 GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
+Table 3: Open loop planning results of imitation-based method, sampling-based method and GAD.
+4.3 Ablation of Ego Feature
+A potentially misleading aspect of current data-driven methodologies in motion planning may be the excessive reliance in utilization of ego history.
+Historical way points are pivotal in predictive tasks for inferring the future trajectories of agents.
+Conversely, planning tasks diverge in that the trajectory is synthesized to achieve specific longitudinal and lateral objectives such as velocity and lateral position, thereby necessitating minimal dependence on historical motion.
+We conduct experiments to validate the necessity of ego history and target speed encoding in Tab.
 2.
-It is noteworthy that the imitation strategy incurs penalties proportional to the trajectory’s divergence from the empirical ground truth, resulting in minimal $L2$ errors across all examined horizons.
-4.3 Closed-loop Experimental
+The historical information is derived from two distinct sources within our framework: one pertaining to rasterized ego motion and the other to vectorized historical positions.
+Since lateral targets can be deduced from route and landmarks, we only provides feature encoding in target speed of 3s-th to guide trajectory generation(it can be conceptualized as a data-driven sampler analogous to the cruise scenario strategy within lattice framework but only concerns the speed in 3s-th).
+Such setting provides a tunable parameter in deployment for the users to adjust the desired cruise speed range.
+As show in the Tab.
+2, we evaluate the open-loop metric and the closed-loop scene passing rate of lane keeping and turning for imitation planner.
+The results indicate that history feature dose help to improve the open-loop metric, but target speed feature is the key for success closed-loop.
+Method without history encoding failed in turning scene.
+4.4 Closed-loop Experiments
 Initially, we conducted an ablation study to assess the necessity of various samplers within a closed road environment.
 Subsequently, we identified the optimal sampler combination for further evaluation on open urban roadways.
-The outcomes of this ablation study are delineated in the appendix.
 Comprehensive real-world testing of these methodologies was conducted in densely populated urban settings, under the vigilant supervision of human safety drivers.
-The specific driving route utilized for our experiments is depicted in the appendix.
+The specific driving route utilized for our experiments is depicted in Sec.
+A.15.
 Throughout the extensive public road testing, spanning over 30 kilometers, the model adeptly executed a diverse array of complex maneuvers, such as lane-following, merging, yielding to pedestrians.
 The results of these road tests are documented in Tab.
-3
+4.
+Method comfort event in 30Km efficiency MPI Lat.
+accel.
+Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute UniAD[23] 0.351 3.12 3.68 - - - - 15.20 0.31 ST-P3[22] 0.857 4.54 8.53 4 3 14 3 8.62 1.36 GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72 GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75 GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75 GAD-gan 0.511 1.15 3.52 3 1 2 1 5.51 4.28 GAD-gan† 0.492 1.16 2.77 0 1 2 1 5.54 7.50 DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
+Table 4: Road test data were systematically collected during a 30 km road test, encompassing the computation of average time per loop.
+Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded.
+(†) indicates boosted by safety layer and cross frame align.
+The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
 It is noteworthy that methodologies exhibiting superior imitation metrics do not invariably demonstrate enhanced performance in closed-loop evaluations.
-This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks.
+This discrepancy may be attributable to the limitations of the imitation metric in adequately assessing the complexities of planning tasks(we illustrated the detail in Sec.
+A.12).
 Nonetheless, the multi-modal imitation generator and GAN-based generator significantly enhance the system’s performance in both open-loop and closed-loop settings, primarily in terms of the planner’s intelligence and the continuity between consecutive frames.
@@
 Closed-loop results suggest that multimodality mitigates the brittleness inherent in imitation learning, while the lattice sampler further enhances the robustness of the entire system.
+See Sec.
+A.7
 To rigorously evaluate the efficacy of the proposed methodology as a substitute for off-the-shelf planners, we conducted a comparative analysis between the GAD and a conventional modularized, user-level, rule-based system, referred to as DPQP in Tab.
-3.
+4.
 Notwithstanding, a discernible performance discrepancy persists in terms of system’s robustness and comfortableness.
 We posit that increase training data size could ameliorate this disparity.
+The system obtains compariable results with DPQP after boosted by the safety layer and cross frame align as mentioned in Sec.
+3.7.
 5 Conclusions
-We introduce a sophisticated DL pipeline from prediction to motion planning in HD map-free setting, augmented with an enhanced safety assurance.
-Our investigations have yielded substantial enhancements over prevailing data-driven methodologies within the complex urban environments of Shanghai.
-This paradigm enables the safety deployment of learned planners in real-world scenarios, offering interpretability and leveraging their capacity to evolve with accumulating data, thereby managing scenarios of greater complexity than those addressed by traditional rule-based systems.
-Future work may integrate generative learning techniques such as Generative Adversarial Networks (GAN), Conditional Variational Autoencoders (CVAE), and diffusion models to further augment the data-driven capabilities in motion planning.
-Furthermore, evidence suggests that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners.
-It is recommended that subsequent research in motion planning incorporate the imitation metric to gauge the intelligence domain of planners and to develop additional metrics for closed-loop evaluation.
+We introduce a sophisticated DL pipeline from prediction to motion planning in an HD map-free setting, augmented with enhanced safety assurance, which has shown substantial enhancements over prevailing data-driven methodologies in Shanghai’s complex urban environments.
+This paradigm enables the safe deployment of learned planners in real-world scenarios, offers interpretability, and leverages their capacity to evolve with accumulating data, managing more complex scenarios than traditional rule-based systems.
+Future work may integrate generative learning techniques such as CVAE[27] and diffusion models[40] to further augment these capabilities.
+Additionally, our results suggest that while imitation learning can enhance motion planning tasks, the imitation metric alone is insufficient as a definitive criterion for assessing the efficacy of motion planners; thus, it is recommended that subsequent research incorporate this metric to gauge the intelligence domain of systems and develop additional metrics for closed-loop evaluation.
 6 Acknowledgements
@@
 A.1 Perception Outputs
-At each timestamp, measured states for each tracked entity are provided in terms of positions, velocities, and accelerations.
-Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead.
+At each timestamp, we have 3D detection results for surround view camera, these inputs are fed into an existing tracking system, outputting stat estimates S for all surrounding actors (state encompasses the bounding box dimensions, spatial coordinates, linear velocity, linear acceleration, directional orientation, and the rate of change of heading).
+Denote a set of discrete times at which tracker outputs state estimates as $[t_{1},t_{2},...t_{T}]$, where time gap between consecutive time steps is constant (the gap is equal to 0.1s for tracker running at a frequency of 10Hz).
+Furthermore, the perception module also provides information about all perceivable landmarks and road boundaries ahead, measured by the polynomial curve.
 Stop lines are also essential when the ADV approaches the junction area.
-The above obstacles and landmarks are either described by coordinates or measured by polynomial curves, referred to as the parametric form.
-The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status, characterized as non-parametric.
+The scene occupancy, on the other hand, is constructed using a grid map with a resolution of 0.2m, where each grid cell is labeled according to its motion status(static or movable).
+We elucidate the rationale behind our architectural that segregates perception from the comprehensive end-to-end stack: Firstly, for mass production friendly concern, the modularized perception provides explaninablity and tunablitiy.
+Secondly, information articulated in parametric form facilitates reasoning at the agent level, not solely for application within the safety layer and the analysis of explainable cases, but also to enhance our trajectory generation processes.
+Thirdly, parameterization encompasses post-tracking and fusion for the results of single or multi-modal detection on raw sensor data which remain formidable unresolved issues, particularly when considering the substitution of these processes with a model to attain comparable results.
+We leave perception involved version to follow up works when those issue resolved in the future.
 A.2 Scene Rasterization
-The state of landmarks and road-bound is rasterized in the same channels, facilitating traffic flow reasoning at intersections.
-To capture the past motion of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers.
-Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the fading effect.
-To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as it’s fading history.
+The primary function of scene rasterization is to map scene information to different channels, enabling the neural network to extract scene semantic information more effectively.
+For example, obstacles, routes, and lane lines are mapped into the RGB channels, respectively.
+Specifically, the states of landmarks and road-boundary are rasterized in the same channel, facilitating traffic flow reasoning at intersections.
+To capture the past motions of all traffic actors, their bounding boxes at consecutive time steps are rasterized on top of map vector layers.
+Each historical actor polygon is rasterized in the same channel as the current polygon but with a reduced brightness level, resulting in the diminishing visual impact over time.
+Brightness level at $t$ is equal to $\max[0,(T-t)*\alpha]$, $T$ is timestamp for current frame, where we set $\alpha=0.05$ (the example shown in Fig.
+2).
+To distinct ADV from other agents, we rasterize ego’s polygon in an isolated channel, as well as its fading history.
+The non-moving obstacles, together with occupancy grids of static label, are rasterized in a different channel since they share similar semantic meaning.
 The rasterized map contains a rendering of traffic light permissibility in junctions, but we tackle it a little bit trick to handle the robustness.
-we render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static agent’s polygone that exhibit permissibilit of current ego route.
-All in all, we obtain a 3D tensor of size (Ha , Wa ,C), with C=5 binary channels for the map.
+We render permitted (green light), yield (unprotected), or prohibited (red light ) by masking stop line with a static polygon to exhibit the permissibility of current ego route.
+The route information is depicted in an isolated channel illustrated in Sec.
+A.6.
+All in all, we obtain a 3D tensor of size $(H,W,C)$, with $H=500,W=500,C=5$ for the map.
 A.3 Instance Vectorization
-Parametric information such as surrounding vehicle positions, lane center lines and lane landmarks, are derived with the form of discretized points to depict the traffic scenes.
-Additionally, route information is obtained to facilitate the ADV in navigating to its intended destination.
-Given that the vehicle’s historical positions are temporally continuous, whereas the lane center lines, lane landmarks, and route information are spatially continuous, they are structured as vectors $\mathbf{v}$ that connect adjacent points.
-Specifically, the $i$th vector can be defined as
-$\mathbf{v}_{i}=[p_{i},p_{i+1},a_{i}]$ (17)
-where $p_{i}$ represents the $i$th point, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
-Similar to the process of scene rasterization, traffic lights are considered as static agents and can be vectorized using unchanged vectors.
+Parametric information for certain instance such as trajectories, lanes, and routes are presented in the form of discrete points.
+In order to enable long-distance interaction, we process these points using a vectorization method to generate a set of instances.
+For each instance $\mathbf{I}_{i}$, it is comprised of a collection of vectors that belong to the same group:
+$\mathbf{I}_{i}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{n}]$ (16)
+As for each vector ${\mathbf{v}_{\textit{i}}}$, it can be defined as:
+$\mathbf{v}_{i}=[\mathbf{p}_{i},\mathbf{p}_{i+1},a_{i}]$ (17)
+where $\mathbf{p}_{i}$ represents the $i$th point coordinates, and $a_{i}$ denotes an attribute that indicates its semantic label such as center lines and visible landmarks.
+Similar to the process of scene rasterization, stop lines are considered as static agents and can be vectorized using unchanged vectors.
 Each vector can be seen as a node in a graph.
 Subsequently, a sequence of nodes belonging to the same category will be combined to form an instance, which will then be interconnected within the graph.
-A.4 Model structure
+A.4 Coordinate
+The above mentioned representations are built upon Ego-ENU coordinate system, rather than the fixed-oriented coordinate system[9], which is derived by rotating the Ego-ENU coordinate by ADV’s heading.
+The Ego-ENU coordinate system is centered around the ADV with the x-axis pointing to the east, the y-axis pointing to the north, and the z-axis pointing upwards.
+In the Ego-ENU coordinate system, the ADV’s historical and future trajectories can be diversely distributed depending on the heading directions, with both x and y coordinates of the same order.
+This allows for a more balanced representation of the vehicle’s motions in both longitudinal and lateral directions.
+Moreover, the fixed-oriented coordinate system tends to extend along the x direction while fluctuating in a small range around 0 along the y direction, leading to challenges in accurately modeling the lateral behaviour such as overtaking and nudging.
+The scale differences between x and y directions causes lateral errors disproportionately large in comparison to longitudinal behaviours.
+Here we represent our comparative experiments to illustrate it.
+We have conducted three experiments to test performances in Ego-ENU and the fixed-oriented coordinate system respectively.
+To tackle the unbalanced distribution of trajectories in fixed-oriented one, we present the original and one with feature normalization.
+Coordiante system with normalization L2(m)
+Ego-ENU x 0.39
+Fixed-Oriented x 0.95
+Fixed-Oriented $\surd$ 0.63
+Table 5: Comparative experiments of coordinate systems.
+According to the test results, it is evident that the L2 distance error of the Ego-ENU coordinate system is the lowest since the predicted future trajectory (noted in blue) approximates to the ground truth(noted in white) and closely follows the route(noted in green).
+Meanwhile, the trajectory in the fixed-oriented coordinate system with feature normalization, also exhibits proximity to the route.
+However, its longitudinal points are irregularly and non-uniformly distributed.
+This irregularity is attributed to the scaling of the longitudinal features to the similar order as the lateral ones.
+Conversely, the trajectory in the fixed-oriented coordinate system without any form of normalization demonstrates the poorest performance, as the end of the trajectory fails to align with the route and deviates outwards, as shown in Fig.
+4.
+Figure 4: Visualization of comparative results.
+(a)the predicted trajectory in Ego-ENU closely follows the route.
+(b)trajectory in the fixed-oriented coordinate system deviates outwards.
+(c)trajectory in the fixed-oriented coordinate system with feature normalization exhibits proximity to the route.
+Therefore, the more symmetrical Ego-ENU coordinate system, which better represents the ADV trajectories, allows for better understanding of vehicle behaviours and enables more effective trajectory planning and overall autonomous driving performance.
+A.5 Model structure
 Our model architecture consists of two main branches.
 The first rasterization branch, utilizes four convolutional blocks with varying scales to extract short and long-range features, which are then used to derive the scene embedding.
-Subsequently, the output from the rasterization branch is fed into two identical-structured decoders, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch.
+Subsequently, the output from the rasterization branch is fed into two decoders with similar structures, each containing four sub-blocks of deconvolution layers, allowing for step-by-step upsampling to match the size of the input from the rasterization branch.
 The decoders generate cost volumes and prediction grid maps.
@@
 The resulting features are concatenated with the instance encoding to create the agent encoding, which is subsequently decoded by an MLP to obtain each agent’s trajectory.
-A.5 Route
+A.6 Route
 As indicated previously, the route plays a critical role in guiding ADVs along roadways.
 Our route information comes from two primary channels: the lane center line and historical commuting trajectories.
-The lane center lines, which can be accessed in perception channels, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane, and if it deviates from the center, it will be guided back to avoid hazards from adjacent lanes.
-However, in lane-changing scenarios, vehicles move across lane markers, making the lane center line ineffective in determining optimal positions for lane changes.
-Consequently, historical commuting trajectories are utilized to capture the lane-change starting and ending positions.
-Upon reaching the lane-change starting position or during the lane-change process, the collected trajectory supersedes the center line as the route, providing guidance for the ADV to change lanes.
-Likewise, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections.
-To address this, the starting and ending positions for turning are recorded.
-As the ADV approaches intersections, human-driving trajectories are employed as the route to guide the ADV through the intersections.
-A.6 Curve Based Sampler
-In this appendix, we present the sampling procedure with more details.
+As shown in Fig.
+5.
+Figure 5: Route visualization.
+(a) lane center line is served as route in cruising scenarios.
+(b) is the turning scenarios, where historical commuting trajectories are served as route, and occupancy provides road boundary information.
+The lane center lines, constructed by landmarks which can be accessed in perception outputs, are mainly utilized for cruising scenarios, as the ADV consistently travels along the middle of the lane.
+We construct three reference lines base on lane center points from current, left and right landmarks to support lane keeping and changing.
+However, in turning scenarios, the absence of landmarks without HD-maps results in the ADV lacking appropriate instructions to navigate intersections.
+To address this, the starting and ending positions from commuting trajectory are recorded.
+As the ADV approaches intersections, the history human-driving trajectories are employed as the route to guide the ADV through the intersections.
+Owing to the absence of the road boundary, the scene occupancy has been adopted to delineate the boundary, as illustrated in Fig.
+5
+Note that in the following visulizations, we omit the occupancy channel for brief.
+A.7 Multi-modal Trajectory Generation
+We provide qualitive result of single modal, multi-modal and GAD-based generator in Fig. 6 and Fig. 7. GAN-based generator help boosting the multi modality in trajectory generation.
+Figure 6: Lane changing scenarios with multi trajectory generation. (a)single trajectory; (b)multi trajectories imitation; (3)GAN-based generator
+Figure 7: Right turning scenarios with multi trajectory generation. (a)single trajectory; (b)multi trajectories imitation; (3)GAN-based generator
+A.8 Curve-based Sampler
 We follow [51] with little modifications for comfort concerns.
 Vehicle trajectories can be categorized into two primary types: straight and turning.
-And turning trajectories encompass lane changes and turning at intersections.
-Lane-changing trajectories are characterized by variable curvature over time, while turning at intersections follows circular-like curves.
+Turning trajectories encompass lane changes and turning at intersections.
+Lane-changing trajectories are characterized by variable curvatures over time, while turning at intersections follows circular-like curves.
 As straight lines can be treated as a special case of curves, we introduce the curve-based sampler that focuses on generating straight, constant curvature circle-like, and variable curvature clothoid-like trajectories.
@@
 Where $\mathbf{s(\xi)}$ is the clothoid curve, $\mathbf{T_{0}}$ and $\mathbf{N_{0}}$ are the tangent and normal vector of point $\mathbf{s_{0}}$, $S(\xi)$ and $C(\xi)$ are associated coefficients and defined as the Fresnel integral.
-As the shapes of clothoid curves solely determined by the scaling factor a, we can generate a variety of clothoid trajectories by considering a suitable range for a, in combination with the initial curvature obtained from the bicycle model and the speed profiles.
-A.7 Lattice Sampler
-The lattice sampler is a reference line-based method used to acquire multiple candidate trajectories along the reference line, taking into account future movements of surrounding obstacles to avoid potential hazards.
-To begin, this method involves sampling multiple candidate trajectories based on the Frenet Frame, where the smooth lane central line is projected to the S coordinate and its normal direction is the L coordinate.
-Trajectories are composed of longitudinal and lateral paths, with the longitudinal one described as the station and time, denoted as $S=\left\{s(t)\right\}$, and the lateral path described as the lateral station and longitudinal one denoted, denoted $L={l(s)}$.
-When sampling trajectories, given the initial state $[s_{0},s_{0}^{{}^{\prime}},s_{0}^{{}^{\prime\prime}}],[l_{0},l_{0}^{{}^{\prime}},l_{0}^{{}^{\prime\prime}}]$ at time $t_{0}$, the end states $[s_{1},s_{1}^{{}^{\prime}},s_{1}^{{}^{\prime\prime}}],[l_{1},l_{1}^{{}^{\prime}},l_{1}^{{}^{\prime\prime}}]$ at time $t_{1}$ are sampled to cover cruising, following, overtaking and stopping maneuvers, and consequently fits polynomials.
-As for cruising maneuvers, vehicles must adhere to a specific cruising speed by adjusting their acceleration and deceleration.
-Therefore a set of prediction time horizon $t_{1}$ and end-condition longitudinal velocities $s_{1}^{{}^{\prime}}$ are uniformly sampled within a pre-defined range of the cruising speed.
-This sampling process is facilitated by a quartic polynomial.
-When it comes to stopping maneuvers, vehicles are obliged to come to a complete stop within a designated distance in order to comply with traffic signals or ensure a safe distance from stationary obstacles.
-In this scenario, the end states such as speed and accelerations are fixed at 0, while the prediction time $t_{1}$ and end longitudinal distance $s_{1}$ are sampled.
-These parameters are then described using a quintic polynomial.
-Regarding following and overtaking maneuvers, these actions often arise when a preceding vehicle is moving at a slower pace or a new vehicle unexpectedly merges into a lane.
-The execution of these complex maneuvers is facilitated through the use of an s-t graph, which depicts the occupation of the lane by vehicles over time as illustrated in Figure x.
-By analyzing the s-t graph, end states can be sampled within unoccupied areas, marked in red (for overtaking) and blue (for following), thus facilitating the completion of these maneuvers through the use of a quintic polynomial.
-Finally, these polynomials can be transformed back to Cartesian coordinate system to obtain the corresponding trajectories, which are consequently evaluated for compliance with kinematic constraints.
-If any trajectory does not meet the requirements, it will be removed from the sampling set.
-A.8 Retrieval-based Sampler
+Given the initial curvature obtained from the bicycle model and the speed profiles, we can generate a variety of clothoid trajectories by considering a suitable range for the scaling factor $a$.
+A.9 Retrieval-based Sampler
 The retrieval-based sampler does not depend on the map or lane information to generate trajectories, instead, it retrieves a cluster of trajectories whose initial state are close to the ADV’s current state from a pre-established expert trajectory dataset.
-To build the expert trajectory dataset, we utilized about 350 hours of driving trajectories from experienced human drivers.
+To build the expert trajectory dataset, we utilized about 200,000 driving trajectories from experienced human drivers.
 However, a substantial portion of the original human driving trajectories contains duplicate or very similar samples.
@@
 In this paper, we assign different weights to the differences in velocity v, acceleration a and curvature $\kappa$, with $\beta_{1}=1.0$, $\beta_{2}=5.0$, $\beta_{3}=40.0$ and all nodes with $D<1.0$ will be added into output for the retrieval-based sampler.
-A.9 Hardware
+A.10 Lattice Sampler
+The lattice sampler is a reference line-based method used to acquire multiple candidate trajectories along the reference line, taking into account future movements of surrounding obstacles to avoid potential hazards.
+To begin, this method involves sampling multiple candidate trajectories based on the Frenet frame, where the smooth lane central line is projected to the S coordinate and its normal direction to the L coordinate.
+Trajectories are composed of longitudinal and lateral paths, with the longitudinal one described as the station and time, denoted as $S=\left\{s(t)\right\}$, and the lateral path described as the lateral station and longitudinal station, denoted as $L={l(s)}$.
+When sampling trajectories, given the initial state $[s_{0},s_{0}^{{}^{\prime}},s_{0}^{{}^{\prime\prime}}],[l_{0},l_{0}^{{}^{\prime}},l_{0}^{{}^{\prime\prime}}]$ at time $t_{0}$, the end states $[s_{1},s_{1}^{{}^{\prime}},s_{1}^{{}^{\prime\prime}}],[l_{1},l_{1}^{{}^{\prime}},l_{1}^{{}^{\prime\prime}}]$ at time $t_{1}$ are sampled to cover cruising, following, overtaking and stopping maneuvers, and consequently fits polynomials.
+Figure 8: Sampling in the Frenet frame, and then converted back to the Cartesian coordinate system.
+As for cruising maneuvers, vehicles must adhere to a specific cruising speed by adjusting their acceleration and deceleration.
+Therefore a set of prediction time horizon $t_{1}$ and end-condition longitudinal velocities $s_{1}^{{}^{\prime}}$ are uniformly sampled within a pre-defined range of the cruising speed.
+This sampling process is facilitated by a quartic polynomial.
+When it comes to stopping maneuvers, vehicles are obliged to come to a complete stop within a designated distance in order to comply with traffic signals or ensure a safe distance from stationary obstacles.
+In this scenario, the end states such as speed and accelerations are fixed at 0, while the prediction time $t_{1}$ and end longitudinal distance $s_{1}$ are sampled.
+These parameters are then described using a quintic polynomial.
+Regarding following and overtaking maneuvers, these actions often arise when a preceding vehicle is moving at a slower pace or a new vehicle unexpectedly merges into a lane.
+The execution of these complex maneuvers is facilitated through the use of an s-t graph, which depicts the occupation of the lane by vehicles over time as illustrated in Fig.
+9.
+By analyzing the s-t graph, end states can be sampled within unoccupied areas, marked in red (for overtaking) and blue (for following), thus facilitating the completion of these maneuvers through the use of a quintic polynomial.
+Figure 9: ST graphs. (a) An obstacle prepares to cut in the current lane, so ADV chooses to overtake; (b) A preceding vehicle moves slowly, so the ADV slow down.
+Finally, these polynomials can be transformed back to Cartesian coordinate system to obtain the corresponding trajectories, which are consequently evaluated for compliance with kinematic constraints.
+If any trajectory does not meet the requirements, it will be removed from the sampling set.
+A.11 Deduction of GAN-based Generator
+Given the optimization target for evaluator:
+$L_{evaluator}=\max_{\hat{s}}[E(\tilde{s})-E(\hat{s})+Diff(\tilde{s},\hat{s})]$ (18)
+The target for generator become:
+$L_{generator}=E(\hat{s})-E(\tilde{s})+Diff(\tilde{s},\hat{s})$ (19)
+$E(\tilde{s})$ is determinate in the generator’s step, thus omitted, and we expand $Diff(\tilde{s},\hat{s})$ as defined in Eq. 6 , Eq. 19 become:
+$L_{generator}=E(\hat{s})+L2(\tilde{s},\hat{s})+O(\hat{s})+G_{rid}(\hat{s})$ (20)
+Where:
+$E(\hat{s})=\sum C(\hat{s})+\alpha O_{cc}(\hat{s})+\beta G_{rid}(\hat{s})$ (21)
+We then merge $O(\hat{s})$ and $G_{rid}(\hat{s})$ into Eq. 21, tuned by hyper-parameter $\alpha$ and $\beta$, thus Eq. 20 become:
+$L_{generator}=E(\hat{s})+L2(\tilde{s},\hat{s})$ (22)
+Here the $L2$ is replaced by multi-path imitation defined in Eq. 10 to ensure diversity, the $\hat{s}$ here is given by the generator in Eq. 15, thereby:
+$L_{generator}=E(G(m))+L_{imit\_plan}(\tilde{s},G(m))$ (23)
+A.12 Limitation of Open-loop Evaluation
+We illustrate the limitation of open-loop evaluation.
+Firstly, a closed-loop planner will adjust to the target in the loop to keep safe and comfortable since real-world planning needs handle the deviations brought by the external disturbances and controller’s modeling errors, which is invisible and omitted, thus cannot be tested by the open-loop planner.
+Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller.
+The open-loop tests will come to an end once the trajectories are generated and compared with the recorded ones.
+However, unexpected and unreasonable jerks and curvatures, which can only be tested when the controller is in the loop, are usually ignored; therefore, the results become less convincing.
+Existing imitation-based approaches formulate the modeling in planning task as prediction problem, which takes ego motion and scene feature to predict future movement.
+Predominantly, these models are assessed using imitation metrics and subjected to open-loop testing.
+Nevertheless, a fundamental discrepancy remains, as the aim of planning transcends mere imitation.
+It is imperative to meticulously craft the learning objectives to avert the entrapment of the model within a local optimum that, while demonstrating proficient motion prediction capacities, fails to adequately fulfill the planning objectives.
+A.13 Hardware
 Our autonomous driving algorithms have undergone rigorous testing on the recently launched Jiyue 01 vehicle, whose detailed parameters can be found in the following link: https://www.jiyue-auto.com/robocar-01.
@@
 More NVIDIA-Orin parameters can be found in the following link: https://developer.nvidia.com/drive/agx.
-A.10 Task Perf
+A.14 Task Perf
 Figure 10: Tasks on board
 Task Time Cost(ms)
-Ave 90p Max
+Ave 90th Max
 Build Frame 0.77 1.13 5.47
@@
 We enumerate the time expenditures associated with each task in Tab.
-4.
+6.
 These tasks are orchestrated by a flow-based scheduler, facilitating concurrent execution as depicted in Fig.
-4.
+10.
 The deployment of the model via TensorRT is conducted without the utilization of acceleration techniques such as quantization or pruning.
 Figure 11: Test Drive Route
-A.11 Road Test
+A.15 Road Test
 Our road tests are conducted in the Jiading District of Shanghai, China, known as ”Shanghai Automobile City”.
@@
 Finally, the route circles back to Yecheng Road and returns to destination No.1688.
-The total length of the test route is 3.0 kilometers long.
-A.12 Dateset
-We collected the record files in 15 hours, which captured the perception output, the ADV trajectory, and route info.
-All the records are post-processed to obtain 200k clips in total.
-By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future.
-We partition the dataset into 16w for training and 4w for testing.
+The total length of the test route is about 3.0 kilometers long.
+The video for road test is available at https://youtu.be/DYPB5lCRvn0
+A.16 Dataset
 We collected 15 hours of record files, containing messages such as perception outputs, ADV trajectories, traffic lights and route information.
@@
 We partition the dataset into 160k for training and 40k for testing.
-A.13 Training details
+A.17 Training details
 The Adam optimizer is used with an initial learning rate of $5e^{-4}$ and a batch size of 32.
@@
 Experiments are conducted with 4 NVIDIA A800 GPUs with 60 epochs to complete.
-In the first stage of our proposed two-stage training model, we regress a grid map by a cnn encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future.
+In the first stage of our proposed two-stage training model, we regress a grid map and cost volume by a CNN encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future.
 Specifically, CNN-encoder embeds the scene image into feature map with 12 convolution layers, and then 4 deconvolution layers are used to decode feature map to the output map.
 In the second stage of obstacle trajectory prediction, we reuse the CNN encoding layer from the first stage and initialize the encoder parameters with the last training step.
-As for obstacle trajectory prediction, we use an LSTM layer to replace the cnn decoder.
+As for obstacle trajectory prediction, we use an LSTM layer to replace the CNN-decoder.
 In order to concentrate more on obstacles, we crop the encoding feature map centered agents, as the local scene feature of obstacles.
@@
 Then, the local scene vector will be sent to the LSTM decoder, concatenated with agent’s history vectorized embedding vector.
+The training methodology for the GAN-based planner follow similar two-stage procedural, undergoing fine-tuning on the pre-trained imitation planner to stabilize and accelerate the training process.
+A.18 Qualitative Results
+We show typical qualitative results in Fig.
+12 - Fig.
+16 to illustrate that our work ensures the ADV can navigate through diverse traffic scenarios to reach the destination safely.
+Note that (a) shows the traffic scenario, and the blue points are the final trajectory selected by our model.
+(b) is the visualization of the cost volume which explains the areas with higher probabilities to travel.
+(c) is the visualization of the grid map to illustrate the future occupancies of obstacles.
+It should be noted that, figure (b) shows the superposition of our predictions for the occupancy at the next 1s, 2s, and 3s.
+The red area represents the highest probability of occupancy for the next 1s for the ego vehicle.
+The green and blue areas represent the highest probability of occupancy for the next 2s and 3s, respectively.
+Figure (c) illustrates the superposition of occupancy at the current moment, in 1s, and in 2s.
+The red region denotes the current position of the obstacle, while the green indicates the areas that the obstacle may occupy at 1s, and the blue represents the areas the obstacle may occupy at 2s.
+Figure 12: Qualitative results: as the vehicle in the right lane accelerates and merges into the ego lane, the ADV opts to decelerate in order to prevent potential rear-end collisions.
+Figure 13: Qualitative results: the ADV is transitioning to the rightmost lane, while also considering the option of using the middle lane if the rightmost lane is occupied.
+Figure 14: Qualitative results: the ADV is maneuvering out of the traffic congestion by overtaking the preceding vehicle at a low speed in order to mitigate potential hazards.
+Figure 15: Qualitative results: the ADV is performing a right turn at a decreased speed to prioritize safety.
+Figure 16: Qualitative results: the ADV is making a left turn at an unprotected intersection. As a vehicle from the oncoming direction approaches, the ADV reduces its speed.